### v1

# jina-clip-v2: Multilingual Multimodal Embeddings for Text and Images

Abstract
Contrastive Language-Image Pretraining (CLIP) is a highly effective method for aligning images and texts in a shared embedding space. These models are widely used for tasks such as cross-modal information retrieval and multi-modal understanding. However, CLIP models often struggle with text-only tasks, underperforming compared to specialized text models. This performance disparity forces retrieval systems to rely on separate models for text-only and multi-modal tasks. In this work, we build upon our previous model, jina-clip-v1, by introducing a refined framework that utilizes multi-task, multi-stage contrastive learning across multiple languages, coupled with an improved training recipe to enhance text-only retrieval. The resulting model, jina-clip-v2, outperforms its predecessor on text-only and multimodal tasks, while adding multilingual support, better understanding of complex visual documents and efficiency gains thanks to Matryoshka Representation Learning and vector truncation. The model performs comparably to the state-of-the-art in both multilingual-multimodal and multilingual text retrieval benchmarks, addressing the challenge of unifying text-only and multi-modal retrieval systems.
1 Introduction
Contrastive text-image pre-training has become a widely adopted approach to building robust text-image alignment models. These models are particularly effective for tasks like cross-modal retrieval and zero-shot classification (Radford et al., 2021), demonstrating strong generalization capabilities. Conventional text-image pre-training excels at aligning text and image embeddings but struggles with aligning text-to-text embeddings (Koukounas et al., 2024). This is because cross-modal retrieval training is inadequate for text-only retrieval. The texts available are typically image captions, which are generally short and information-poor, and training procedures don’t use standard techniques for training text embeddings, like hard negatives. (Zhang et al., 2024a)
To overcome these limitations, Koukounas et al. (2024) introduced a multi-task, multi-stage contrastive learning approach to simultaneously align text-text and text-image embeddings. Their method involves three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2) and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3). The resulting model, jina-clip-v1, achieves strong performance on both the cross-modal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and text embedding MTEB Benchmark (Muennighoff et al., 2023).
Despite its strengths, this model has several limitations. First, it is an English-only multimodal embedding model, which makes it unsuitable for multilingual document retrieval. Second, jina-clip-v1 struggles with images that contain text, tables, graphs, diagrams, and other complex content (Faysse et al., 2024).
To address these challenges, we propose an enhanced text-image pre-training scheme. Our approach incorporates multilingual text-image and text-text pairs, along with text-image pairs containing text and other difficult-to-process complex visual structures. Additionally, we gradually increase image resolution during pre-training to balance computational efficiency and model performance. The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on cross-modal retrieval benchmarks against state-of-the-art models like NLLB-CLIP-SigLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the Retrieval and STS tasks of the multilingual MTEB benchmark (Muennighoff et al., 2023). Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on benchmarks for visually rich document retrieval, compared to jina-clip-v1. To reduce vector storage costs, we employ Matryoshka Representation Learning (Kusupati et al., 2024) to enable dimension truncation on the output vectors with minimal performance degradation.
2 Related Work
Radford et al. (2021) pioneered the dual encoder architecture trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pre-trained vision-language models. Building upon this foundation, Zhai et al. (2023) propose an alternative and more efficient sigmoid objective, while Sun et al. (2023) optimize the model’s dimensions and scale up both dataset and model sizes, achieving state-of-the-art performance on cross-modal tasks. Koukounas et al. (2024) enhance a CLIP-based retriever with strong text-to-text and cross-modal retrieval capabilities, though fall short in addressing retrieval scenarios that involve candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022). To mitigate the modality bias, Lin et al. (2024) take a different approach by fine-tuning a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks.
Extensive research has also focused on extending CLIP’s capabilities to multilingual contexts. To address the lack of image-caption pairs in languages other than English, Carlsson et al. (2022) apply knowledge distillation to retrain the text encoder using machine translation data. NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
The importance of good information retrieval in Retrieval-Augmented Generation (RAG) systems has motivated rapid progress in multilingual text embedding models. These models are typically based on either a robust encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020) or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023). Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, leveraging extensive multilingual training with instruction tuning and multi-task learning techniques, respectively. mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the context length. Similarly, Sturua et al. (2024) employ RoPE to increase the model’s context length and apply LoRA tuning (Hu et al., 2021) to optimize embeddings for downstream tasks.
3 Training
| Feature | Text Encoder | Image Encoder |
| Base Model | Jina-XLM-RoBERTa (Sturua et al., 2024) | EVA02 L/14 (Sun et al., 2023) |
| Parameters | 561M | 304M |
| Input Specification | 8,192 tokens (max) | 512×512 pixels |
| Output Dimensions | 1,024 | 1,024 |
| Layers | 24 | 24 |
| Attention Implementation | FlashAttention2 (Dao et al., 2022) | xFormers (Lefaudeux et al., 2022) |
| Pooling Strategy | Mean pooling | CLS pooling |
| Additional Features | 89 languages supported | Patch size 14x14 |
The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024). One encoder is responsible for encoding texts in a 1024-dimensional embedding space, while a second encoder embeds images in the same embedding space.
The text encoder is initialized with the pre-trained Jina-XLM-RoBERTa model weights. Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the XLM-RoBERTa (Conneau et al., 2020) weights to a modern encoder-only architecture that adds Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and LoRA (Hu et al., 2021).
For the image encoder, we have opted for the EVA02 (Fang et al., 2023b) family of ViT models, similarly to jina-clip-v1. We have selected a relatively large pre-trained model, similar in size to the text encoder. This model implementation includes 2D rotary positional embeddings (Su et al., 2023) and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022).
Architectural details for both encoders are presented in Table 1.
3.1 Datasets
Similarly to jina-clip-v1 (Koukounas et al., 2024), we construct four different training datasets. Dataset denotes a dataset of text pairs, denotes a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset. All datasets are multi-lingual.
The text corpora and are introduced in Sturua et al. (2024) and contain data in 30 languages. The text pair corpus consists of query and target pairs from a diverse collection of text-pair datasets. The triplet text corpus is a high-quality dataset with hard negatives from various sources. Each training item contains one annotated positive and seven negative items. Regarding , we focus on Retrieval and Semantic Textual Similarity (STS) datasets for the scope of this work.
Our multi-modal training data comes from multiple sources. For the English part, we rely on the DFN dataset (Fang et al., 2023a), a collection of 400M image-caption pairs, filtered using data filtering networks. For the rest of the languages, we collect 400M samples from CommonPool (Gadre et al., 2023), using multilingual CLIP to identify and discard low-quality items. All images are resized to . These two large-scale datasets are included in and a small part with captions longer than 256 tokens is held out for .
Inspired by visual QA applications (Faysse et al., 2024), We have diversified our training data with PDFs, scientific graphs, infographics and Wikipedia images. The following datasets are included in both and : DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024). For the QA datasets, to obtain a caption for an image, we concatenate query and answer for each sample. For the WIT dataset, we use the reference caption in and a concatenation of caption, page title, section tile, page description and section description for .
Finally, includes multilingual synthetic long captions, similar to how Koukounas et al. (2024) make use of ShareGPT4v (Chen et al., 2023a). We use the GPT4v API (OpenAI, 2023) to generate detailed image descriptions for 40,000 images in 30 languages. This adds 1.2M generated multilingual long captions to .
3.2 Training Stages
Following the strategy proposed in Koukounas et al. (2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
Stage 1 focuses on aligning the multi-modal representations while also improving the text representations of the text encoder. We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes. Towards the end, when performance peaks, we switch to a higher image resolution of using positional embedding interpolation as a warm-up for the next stage.
Stage 2 trains on and , with context length increased from 77 to 512 and image resolution set at . This stage improves text embedding performance on longer text lengths, while maintaining alignment between the two modalities.
Stage 3 uses hard negatives from to further improve the text encoder in distinguishing relevant from irrelevant text. To maintain text-image alignment, we continue training with . We do one more positional embedding interpolation to increase the image resolution to and train on this resolution for the duration of stage 3.
Appendix table 6 specifies the training details for each step.
3.3 Loss Functions
Given a batch of embedding pairs , the InfoNCE loss (Van den Oord et al., 2018), given in equation 1, evaluates the cosine similarity between query and its corresponding target , relative to the similarity of all other targets in the batch. The loss is calculated in both directions to preserve the symmetry of similarity measures. The temperature parameter influences how the loss function weighs minor differences in the similarity scores.
When hard negatives are available for query , an extended version of the loss is used. Given a batch of samples , consisting of a query , a positive match , and seven negatives , the extended loss function, denoted here as , is given in equation 2. Similarly to , this loss function is bidirectional.
Equation 3 formulates the loss for each of the three stages. In each stage, we optimize a joint loss function, i.e. the sum of two loss functions, one operating on text representations and one on multi-modal representations. All stages optimize except Stage 3, which uses hard negatives in the text loss branch and thus calculates .
In Equation 3, denotes a batch of text pairs drawn from the dataset of text pairs , while denotes a batch of text samples with negatives drawn from the hard negatives dataset . Similarly denotes a batch of text-image pairs drawn from the multimodal short-caption dataset , while denotes a batch of text-image pairs sampled from the multimodal long-caption dataset .
3.4 Matryoshka Representation Learning
In all stages of training, each loss function is evaluated on the full 1024-dimension output vectors of the two encoders as well as on truncated vectors of 64, 128, 256, 512 and 768 dimensions. In other words, each loss component on equation 3 is the sum of the same loss function evaluated at different embedding dimensionalities.
This training technique, called Matryoshka Representation Learning, (Kusupati et al., 2024) trains the model to learn the most important features in the vector sequence. This makes it possible, at inference time, to truncate embedding vectors, thereby reducing storage and computation costs when comparing vectors, at a minor performance penalty.
4 Evaluation
We have evaluated the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual cross-modal retrieval benchmarks in Section 4.1, common text embedding tasks from the MTEB benchmark suite in Section 4.2, visual document retrieval tasks in Section 4.3, and an ablation study of jina-clip-v2’s matryoshka representations in Section 4.4.
Detailed evaluation results are presented in the appendices. Detailed CLIP Benchmark results are given in Appendix tables 7, 8 and 9, MTEB results are in in Appendix 10. ViDoRe benchmark results are in Appendix 11. A full MRL evaluation can be found in tables 12, 13, 14, 15 in the appendices.
4.1 Cross-modal Evaluation
| Benchmark | CLIP Benchmark - EN | Crossmodal-3600 | Multilingual MS COCO | |||
| Task Type | Zero-Shot Retrieval | |||||
| Model - Metric | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
| nllb-siglip-large | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
| nllb-siglip-base | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
| jina-clip-v1 | 77.75 | 87.65 | - | - | - | - |
| jina-clip-v2stage 1 | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
| jina-clip-v2stage 2 | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
| jina-clip-v2 | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |
T-I r@5 : Text to Image Recall@5 [%] I-T r@5 : Image to Text Recall@5 [%]
We evaluate jina-clip-v2 on a set of English and multilingual cross-modal tasks, conducting a comparative analysis against both its predecessor jina-clip-v1 and the NLLB-CLIP-SigLIP (Visheratin, 2023b) large and base variants, a top-performing multilingual CLIP model. Additionally, we track the model’s performance at different training stages to evaluate the effectiveness of different parts of the training protocol. Measuring 865M parameters, our model sits between the two NLLB-CLIP-SigLIP versions in terms of size: nllb-siglip-base (507M parameters, 41% smaller than ours) and nllb-siglip-large (1.2B parameters, 39% larger than ours).
For English zero-shot image-text and text-image retrieval, we conduct evaluations on Flickr30K (Young et al., 2014) and MS COCO Captions (Chen et al., 2015), both of which are included in the CLIP Benchmark. For multilingual cross-modal tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and multilingual MS COCO (Chen et al., 2015) datasets.
As shown in Table 2, jina-clip-v2 demonstrates strong performance across both English and multilingual benchmarks while showing clear progression through its training stages. On the English CLIP Benchmark, the final model achieves 79.09% and 89.73% on text-to-image and image-to-text retrieval respectively, showing significant improvement from stage 1’s 73.87% and 86.61%. While nllb-siglip-large leads in text-to-image retrieval at 81.54%, jina-clip-v2 outperforms its predecessor jina-clip-v1 and surpasses both NLLB-CLIP-SigLIP variants in image-to-text retrieval by 2.1% and 1.6% respectively. On the multilingual front, we obtain competitive results on Crossmodal-3600, achieving 81.43% and 83.23% on text-to-image and image-to-text retrieval respectively, approaching the performance of nllb-siglip-large. On multilingual MS COCO (Chen et al., 2015), jina-clip-v2 performs comparably to NLLB-CLIP-SigLIP models, reaching 84.87% and 86.03% on text-to-image and image-to-text retrieval. The improvement from Stage 1 to the final model is notable across all benchmarks, with gains of up to 10.6 percentage points on Crossmodal-3600’s text-to-image retrieval, demonstrating the effectiveness of the training strategy.
4.2 Text Retrieval and STS Evaluation
| Benchmark | Retrieval | STS | ||
| Metric | nDCG@10 [%] | Spearman Correlation [%] | ||
| Model - Language | EN | Multilingual | EN | Multilingual |
| nllb-siglip-large | 24.91 | - | 74.89 | - |
| jina-embeddings-v3 | 53.87 | 72.58 | 85.80 | 69.81 |
| jina-clip-v1 | 48.33 | - | 80.92 | - |
| jina-clip-v2stage 1 | 40.91 | - | 79.67 | - |
| jina-clip-v2stage 2 | 43.17 | - | 80.33 | - |
| jina-clip-v2 | 49.32 | 69.85 | 81.29 | 67.77 |
Table 3 presents the evaluation on MTEB tasks, demonstrating the effectiveness of jina-clip-v2’s text encoder for both retrieval and semantic textual similarity (STS) tasks. On English retrieval tasks (nDCG@10), jina-clip-v2 achieves an average score of 49.32, a significant improvement over nllb-siglip-large’s 24.91 and a slight improvement over jina-clip-v1’s 48.33, though not reaching jina-embeddings-v3’s score of 53.87. For multilingual retrieval, the model achieves 69.85, approaching jina-embeddings-v3’s performance of 72.58. In STS tasks measured by Spearman Correlation, jina-clip-v2 demonstrates competitive performance with 81.29 for English tasks, surpassing nllb-siglip-large’s 74.89 and comparable to jina-clip-v1’s 80.92, though below jina-embeddings-v3’s 85.8. On multilingual STS tasks, the model achieves 67.77, slightly below jina-embeddings-v3’s 69.8.
4.3 Visual Document Retrieval
| Model | Average (nDCG@5) [%] |
| jina-clip-v1 | 17.72 |
| nllb-siglip-large | 46.55 |
| jina-clip-v2 | 52.65 |
| jina-clip-v2stage 1 | 37.37 |
| jina-clip-v2stage 2 | 47.76 |
Visual document retrieval requires vision models to capture fine-grained information from images to accurately match documents with queries. This task is particularly challenging as it deals with high-resolution document images, which differ significantly from typical image-caption datasets. As shown in table 4, jina-clip-v2 achieves state-of-the-art performance on the ViDoRe Benchmark (Faysse et al., 2024) for visual document understanding, with an average nDCG@5 score of 52.65%, significantly outperforming both its predecessor jina-clip-v1 (17.72%) and nllb-siglip-large (46.55%). The progression through training stages demonstrates substantial improvements, advancing from 37.37% in stage 1 to 47.76% in stage 2, before reaching peak performance in the final model. These results highlight the effectiveness of our training strategy and the impact of the new datasets introduced for document retrieval tasks.
4.4 Matryoshka Representation Learning
| Dataset - Dimension |
| Text-Image Retrieval - Recall@5 [%] | ||||||||||||
| Avg CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
| Avg Crossmodal-3600 | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| Avg Multilingual MS COCO | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
| Image-Text Retrieval - Recall@5 [%] | ||||||||||||
| Avg CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
| Avg Crossmodal-3600 | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| Avg Multilingual MS COCO | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
| Text-Text Retrieval - nDCG@10 [%] | ||||||||||||
| Avg Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
| Semantic Textual Similarity - Spearman Correlation [%] | ||||||||||||
| Avg STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
Table 5 presents the impact of embedding truncation on the performance of our model. The table is divided into 4 task categories: text-to-image retrieval, image-to-text retrieval, text-to-text retrieval and semantic textual similarity (STS).
The results show that performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks. At 256 dimensions, representing a 75% reduction in embedding size, the model effectively preserves essential semantic information. Substantial performance degradation is only observed at very low dimensions (128 and 64), indicating the model’s very robust acquisition of image and text features.
5 Analysis
5.1 The Role of Image Resolution in Complex Document Retrieval
Our empirical study demonstrates that image resolution plays a critical role in retrieving visually rich documents. To better understand the relationship between image resolution and model performance on such documents, we have conducted a targeted experiment.
We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of (224, 224). This model checkpoint was trained on 1.45 billion image-caption pairs and 1.45 billion text-text pairs. We conducted four runs, keeping the image resolution to for the first run and increasing to , , and for three additional runs. Each run is trained for 3500 steps using the same visually rich training set and hyperparameters. The models are then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents, based on text queries.
We have plotted the results in Figure 2. Unsurprisingly, increasing image resolution has a positive impact on linking captions to visually rich documents. The most significant improvement occurs when resolution increases from to , with the average nDCG@5 score across 10 benchmarks rising from to . A further increase in resolution to also brings a noticeable gain in performance.
Considering that the model is trained with a patch size of 14, increasing the resolution beyond significantly raises the number of patches, a increase when moving from to . This leads to substantial additional computational overhead while yielding only marginal improvement, with nDCG@5 increasing by just . Therefore, we identify as an optimal image resolution, offering a good balance between performance and computational efficiency.
5.2 Unified Batch vs Multi-Task Learning
Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multi-modal alignment task for images and text, and second text-only task. Both use the InfoNCE (Van den Oord et al., 2018) objective. Regardless of whether the loss is computed on image-text or text-text pairs, the same loss function is applied to pairs of vectors. This raises the question: Would unifying the two batches of pairs into a single batch, irrespective of modality, and optimizing a single contrastive objective, offer any benefits?
We refer to this approach as the unified batch technique, in contrast to the multi-task learning paradigm. We present the training paradigm in Figure 3. The motivation is two-fold: first, simplicity of computation and secondly, modality gap mitigation (Liang et al., 2022) by using in-batch negatives across modalities. Compared to multi-task training, the unified batch approach utilizes a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space. Additionally, we integrate a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representation.
However, our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to cross-modal tasks. This was reflected in the slow decay of the loss temperature (our temperature is a trainable parameter in this case) and its eventual stabilization around . This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training, leading to better performance on cross-modal tasks.
We hypothesize that this limitation arises from the fundamental information asymmetry between the visual and textual modalities (Schrodi et al., 2024). Images include fine-grained details of information whereas the corresponding textual descriptions include more coarse-grained level of information (Schrodi et al., 2024; Chen et al., 2023b). In temperature-scaled contrastive learning, the temperature parameter controls the hardness of the loss function with respect to the negative samples, i.e., the level of penalties on the hard negative samples (Wang & Liu, 2021). From an information-theory perspective, the level of penalty on the hard negative text-text samples theoretically differs from that of image-text samples due to the information gap problem. Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is potentially theoretically suboptimal.
6 Conclusion
In this work we present an enhanced strategy for training dual-encoder vision-language embedding models for cross-modal and text retrieval tasks, using contrastive learning. We introduce improvements to the pipeline, namely support for multiple languages, Matryoshka Representation Learning, and finer visual perception. We have trained and released a new model, jina-clip-v2, which has strong cross-modal and text retrieval performance on standard benchmarks. Finally, we have conducted analyses of three important considerations for CLIP models going forward: the ability to understand complex visual inputs performance how the image resolution affects it, and the effects of efforts to mitigate the modality gap.
Appendix A Appendix
| Parameter | Stage 1 | Stage 2 | Stage 3 |
| Image encoder weights init | EVA02 ViT L/14 (Sun et al., 2023) | Stage 1 | Stage 2 |
| Text encoder weights init | JinaXLMRoBERTa (Sturua et al., 2024) | Stage 1 | Stage 2 |
| Peak image encoder LR | 2e-4 | 5e-5 | 5e-6 |
| Image encoder layer-wise LR decay | 1 | 0.98 | 1 |
| Peak text encoder LR | 1e-4 | 5e-5 | 1e-4 |
| Text encoder layer-wise LR decay | 1 | 0.98 | 1 |
| Image-text pairs batch size | |||
| Text pairs batch size | |||
| Total steps | |||
| Max sequence length | |||
| Image-text pairs samples seen | 1.7B | 50M | 16M |
| Text pairs samples seen | 1.7B | 50M | 2M |
| Number of GPUs - H100s 80GB | 8 | 8 | 8 |
| LR schedule | cosine decay | ||
| Optimizer | AdamW (Loshchilov & Hutter, 2019) | ||
| Optimizer hyper-parameters | |||
| Weight decay | 0.02 | ||
| Input resolution | (224…384, 224…384) | (384, 384) | (512, 512) |
| Patch size | (14, 14) | ||
| Numerical precision | bfloat16 | ||
| Dataset - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||||
| Flickr30K (Young et al., 2014) | 89.84 | 86.84 | 90.04 | 89.04 | 92.24 | 90.02 | ||||||||||
| MS COCO (Chen et al., 2015) | 68.35 | 60.91 | 69.59 | 66.42 | 70.84 | 69.13 | ||||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||||
| Flickr30K (Young et al., 2014) | 98.00 | 96.10 | 97.40 | 96.40 | 97.10 | 95.00 | ||||||||||
| MS COCO (Chen et al., 2015) | 81.46 | 77.12 | 81.74 | 79.02 | 79.20 | 77.76 | ||||||||||
| Language - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
| average | 81.43 | 73.51 | 84.13 | 82.07 | 79.29 | |||||||||
| ar | 73.56 | 66.22 | 76.89 | 78.92 | 76.94 | |||||||||
| bn | 63.78 | 50.19 | 68.11 | 75.19 | 74.19 | |||||||||
| da | 85.39 | 76.53 | 87.97 | 87.14 | 86.64 | |||||||||
| de | 91.25 | 84.64 | 92.42 | 89.56 | 87.25 | |||||||||
| el | 75.03 | 68.53 | 77.92 | 77.83 | 71.97 | |||||||||
| en | 75.83 | 69.78 | 77.78 | 73.11 | 72.22 | |||||||||
| es | 83.64 | 78.28 | 86.28 | 82.64 | 79.97 | |||||||||
| fi | 82.83 | 75.28 | 85.89 | 86.42 | 81.44 | |||||||||
| fr | 88.78 | 83.50 | 90.67 | 87.86 | 85.58 | |||||||||
| hi | 55.25 | 42.03 | 59.47 | 60.31 | 58.08 | |||||||||
| id | 84.22 | 73.69 | 86.94 | 86.31 | 84.17 | |||||||||
| it | 88.33 | 81.19 | 89.64 | 85.94 | 82.67 | |||||||||
| ja | 87.03 | 77.28 | 90.03 | 86.06 | 83.14 | |||||||||
| ko | 78.81 | 71.81 | 83.22 | 78.75 | 75.47 | |||||||||
| nl | 82.56 | 75.72 | 84.47 | 81.69 | 78.86 | |||||||||
| no | 81.08 | 71.97 | 83.08 | 82.69 | 79.97 | |||||||||
| pl | 84.00 | 79.33 | 86.50 | 82.72 | 78.61 | |||||||||
| pt | 82.42 | 76.17 | 85.19 | 82.69 | 79.44 | |||||||||
| ro | 89.36 | 82.92 | 92.22 | 90.03 | 86.17 | |||||||||
| ru | 88.97 | 82.97 | 91.11 | 86.44 | 83.78 | |||||||||
| sv | 78.06 | 71.33 | 80.56 | 79.33 | 76.17 | |||||||||
| th | 81.61 | 70.92 | 85.08 | 81.14 | 78.83 | |||||||||
| tr | 81.31 | 75.31 | 84.53 | 83.47 | 81.00 | |||||||||
| uk | 88.56 | 82.28 | 89.89 | 85.44 | 81.89 | |||||||||
| vi | 86.64 | 76.31 | 89.06 | 85.56 | 82.56 | |||||||||
| zh | 78.97 | 66.97 | 82.50 | 76.56 | 74.64 | |||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
| average | 83.23 | 79.64 | 84.12 | 80.16 | 76.56 | |||||||||
| ar | 76.25 | 72.14 | 76.67 | 75.86 | 73.69 | |||||||||
| bn | 69.00 | 63.78 | 70.22 | 75.58 | 73.61 | |||||||||
| da | 88.53 | 84.42 | 88.86 | 86.53 | 84.69 | |||||||||
| de | 92.47 | 88.42 | 92.47 | 87.50 | 84.44 | |||||||||
| el | 73.33 | 73.22 | 75.61 | 74.81 | 68.89 | |||||||||
| en | 78.58 | 74.33 | 79.61 | 70.81 | 69.08 | |||||||||
| es | 86.28 | 81.19 | 87.36 | 81.19 | 77.00 | |||||||||
| fi | 84.19 | 81.17 | 85.64 | 84.25 | 78.75 | |||||||||
| fr | 90.89 | 86.56 | 91.00 | 86.64 | 83.64 | |||||||||
| hi | 61.64 | 54.39 | 61.56 | 61.89 | 59.83 | |||||||||
| id | 86.31 | 83.64 | 87.64 | 84.33 | 81.89 | |||||||||
| it | 90.17 | 84.31 | 90.53 | 83.50 | 78.64 | |||||||||
| ja | 88.50 | 85.53 | 89.44 | 84.03 | 81.14 | |||||||||
| ko | 81.42 | 79.22 | 83.00 | 76.75 | 74.17 | |||||||||
| nl | 82.47 | 77.94 | 83.33 | 79.28 | 74.33 | |||||||||
| no | 83.75 | 77.22 | 84.42 | 82.08 | 77.03 | |||||||||
| pl | 84.61 | 83.72 | 85.69 | 79.58 | 75.39 | |||||||||
| pt | 83.94 | 79.31 | 84.06 | 79.39 | 74.89 | |||||||||
| ro | 91.31 | 87.39 | 92.08 | 88.83 | 83.58 | |||||||||
| ru | 90.64 | 87.75 | 90.89 | 84.64 | 80.61 | |||||||||
| sv | 78.28 | 77.53 | 80.11 | 76.58 | 72.31 | |||||||||
| th | 81.94 | 80.22 | 84.11 | 79.56 | 76.28 | |||||||||
| tr | 82.67 | 80.36 | 83.81 | 80.36 | 77.69 | |||||||||
| uk | 89.53 | 86.42 | 89.97 | 83.22 | 78.67 | |||||||||
| vi | 88.06 | 85.64 | 88.58 | 83.14 | 80.03 | |||||||||
| zh | 79.22 | 74.94 | 80.42 | 73.83 | 70.22 | |||||||||
| Language - Model |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
| average | 84.87 | 80.66 | 86.11 | 87.60 | 86.23 | |||||||||
| de | 85.70 | 80.40 | 86.80 | 88.30 | 87.00 | |||||||||
| en | 89.40 | 84.00 | 89.40 | 89.40 | 88.80 | |||||||||
| es | 85.90 | 82.70 | 87.40 | 88.20 | 86.70 | |||||||||
| fr | 85.10 | 80.90 | 87.30 | 87.70 | 86.90 | |||||||||
| it | 85.80 | 83.20 | 87.10 | 89.30 | 87.80 | |||||||||
| ko | 82.10 | 78.30 | 83.00 | 85.20 | 83.60 | |||||||||
| pl | 86.50 | 81.10 | 88.00 | 89.40 | 87.60 | |||||||||
| ru | 81.10 | 79.30 | 82.10 | 83.40 | 82.40 | |||||||||
| tr | 83.70 | 81.30 | 86.30 | 88.30 | 86.80 | |||||||||
| zh | 83.40 | 75.40 | 83.70 | 86.80 | 84.70 | |||||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
| average | 86.03 | 83.02 | 86.45 | 85.37 | 84.56 | |||||||||
| de | 86.20 | 82.90 | 86.90 | 84.40 | 84.30 | |||||||||
| en | 90.50 | 87.80 | 90.60 | 88.30 | 87.50 | |||||||||
| es | 87.00 | 84.90 | 87.50 | 85.80 | 85.20 | |||||||||
| fr | 85.20 | 82.20 | 85.00 | 86.30 | 84.40 | |||||||||
| it | 88.00 | 83.10 | 87.70 | 86.70 | 85.60 | |||||||||
| ko | 82.10 | 79.30 | 84.00 | 83.20 | 82.60 | |||||||||
| pl | 88.50 | 84.50 | 88.90 | 86.80 | 86.60 | |||||||||
| ru | 83.10 | 79.60 | 81.80 | 80.90 | 81.20 | |||||||||
| tr | 85.60 | 83.20 | 85.90 | 87.10 | 85.40 | |||||||||
| zh | 84.10 | 82.70 | 86.20 | 84.20 | 82.80 | |||||||||
| Dataset - Model |
| Retrieval - nDCG@10 | ||||||||||||||||
| Average Retrieval | 49.33 | 53.87 | 40.91 | 43.17 | 24.92 | 48.33 | ||||||||||
| ArguAna | 43.55 | 54.33 | 41.87 | 45.84 | 40.63 | 49.36 | ||||||||||
| ClimateFEVER | 30.12 | 42.36 | 19.32 | 23.74 | 15.57 | 24.81 | ||||||||||
| CQADupstackRetrieval | 41.20 | 42.36 | 41.57 | 42.09 | 21.06 | 40.92 | ||||||||||
| DBPedia | 37.44 | 41.00 | 22.04 | 25.48 | 20.46 | 36.64 | ||||||||||
| FEVER | 84.96 | 89.05 | 60.97 | 66.13 | 14.98 | 76.28 | ||||||||||
| FiQA2018 | 41.93 | 47.35 | 41.87 | 40.97 | 12.90 | 38.27 | ||||||||||
| HotpotQA | 60.14 | 64.67 | 49.28 | 47.00 | 22.38 | 61.89 | ||||||||||
| MSMARCO | 37.47 | 40.82 | 23.06 | 27.05 | 13.35 | 36.91 | ||||||||||
| NFCorpus | 32.89 | 36.62 | 30.95 | 32.11 | 18.97 | 33.52 | ||||||||||
| NQ | 57.17 | 64.23 | 39.78 | 42.80 | 16.96 | 58.09 | ||||||||||
| QuoraRetrieval | 88.14 | 89.09 | 87.12 | 86.92 | 80.71 | 87.88 | ||||||||||
| SCIDOCS | 18.90 | 19.81 | 20.51 | 20.91 | 10.62 | 20.24 | ||||||||||
| SciFact | 65.34 | 72.31 | 68.26 | 67.61 | 35.28 | 67.34 | ||||||||||
| TRECCOVID | 76.73 | 77.72 | 51.29 | 58.63 | 38.15 | 71.61 | ||||||||||
| Touche2020 | 23.94 | 26.30 | 15.72 | 20.21 | 11.75 | 21.15 | ||||||||||
| STS - Spearman correlation based on cosine similarity | ||||||||||||||||
| Average STS | 81.29 | 85.80 | 79.67 | 80.33 | 74.90 | 80.92 | ||||||||||
| BIOSSES | 83.00 | 88.69 | 83.55 | 83.62 | 65.07 | 83.75 | ||||||||||
| SICK-R | 82.38 | 89.62 | 78.96 | 78.63 | 75.83 | 78.95 | ||||||||||
| STS12 | 76.72 | 82.43 | 73.41 | 74.57 | 73.60 | 73.52 | ||||||||||
| STS13 | 79.90 | 89.49 | 80.25 | 82.35 | 76.50 | 83.24 | ||||||||||
| STS14 | 77.49 | 84.95 | 75.13 | 76.89 | 76.35 | 78.68 | ||||||||||
| STS15 | 86.42 | 89.31 | 84.98 | 85.62 | 80.90 | 87.46 | ||||||||||
| STS16 | 85.18 | 86.85 | 83.70 | 82.66 | 74.73 | 83.77 | ||||||||||
| STS17 | 87.87 | 89.99 | 86.13 | 86.95 | 83.84 | 89.77 | ||||||||||
| STS22 | 67.07 | 67.28 | 66.31 | 67.58 | 62.52 | 65.15 | ||||||||||
| STSBenchmark | 86.87 | 89.44 | 84.27 | 84.40 | 79.64 | 84.93 | ||||||||||
| Task |
| Retrieval - nDCG@5 | |||||||||||||
| Average | 52.65 | 37.37 | 47.76 | 17.72 | 46.55 | ||||||||
| ArxivQ | 64.92 | 58.78 | 66.14 | 25.40 | 30.44 | ||||||||
| DocQ | 24.64 | 21.16 | 24.81 | 11.90 | 23.82 | ||||||||
| InfoQ | 57.90 | 56.62 | 56.63 | 35.50 | 59.87 | ||||||||
| TabF | 45.91 | 37.28 | 43.18 | 20.20 | 70.68 | ||||||||
| TATQ | 30.25 | 12.09 | 25.79 | 3.30 | 20.00 | ||||||||
| Shift | 34.07 | 28.41 | 31.54 | 3.80 | 30.79 | ||||||||
| AI | 68.07 | 32.95 | 55.64 | 15.20 | 47.90 | ||||||||
| Energy | 62.15 | 49.15 | 60.97 | 19.70 | 64.94 | ||||||||
| Gov | 68.97 | 37.81 | 55.56 | 21.40 | 58.62 | ||||||||
| Health | 69.05 | 39.50 | 57.37 | 20.80 | 58.43 | ||||||||
| Dataset - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| Flickr30K (Young et al., 2014) | 89.84 | 89.98 | 89.70 | 89.20 | 87.22 | 81.60 | ||||||
| MS COCO (Chen et al., 2015) | 68.35 | 68.26 | 68.16 | 67.43 | 64.58 | 59.41 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| Flickr30K (Young et al., 2014) | 98.00 | 98.10 | 98.00 | 98.00 | 96.30 | 93.40 | ||||||
| MS COCO (Chen et al., 2015) | 81.46 | 81.10 | 81.10 | 80.70 | 78.66 | 73.00 | ||||||
| Language - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| average | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
| ar | 73.56 | 73.39 | 73.17 | 72.61 | 68.42 | 62.28 | ||||||
| bn | 63.78 | 63.67 | 63.64 | 62.39 | 57.58 | 49.58 | ||||||
| da | 85.39 | 85.31 | 84.67 | 84.53 | 81.69 | 75.69 | ||||||
| de | 91.25 | 91.28 | 91.47 | 91.47 | 88.75 | 84.44 | ||||||
| el | 75.03 | 75.08 | 75.25 | 74.69 | 72.00 | 66.81 | ||||||
| en | 75.83 | 75.97 | 76.03 | 75.72 | 74.03 | 69.67 | ||||||
| es | 83.64 | 83.67 | 83.86 | 83.14 | 81.00 | 76.19 | ||||||
| fi | 82.83 | 83.03 | 82.61 | 81.67 | 79.44 | 74.94 | ||||||
| fr | 88.78 | 88.75 | 88.53 | 88.28 | 86.92 | 82.14 | ||||||
| hi | 55.25 | 54.97 | 55.14 | 54.14 | 50.06 | 42.33 | ||||||
| id | 84.22 | 84.00 | 84.11 | 83.25 | 80.00 | 74.17 | ||||||
| it | 88.33 | 88.53 | 88.31 | 87.83 | 85.11 | 79.75 | ||||||
| ja | 87.03 | 87.03 | 86.89 | 86.33 | 82.92 | 75.39 | ||||||
| ko | 78.81 | 78.72 | 78.56 | 77.03 | 73.44 | 65.78 | ||||||
| nl | 82.56 | 82.56 | 82.31 | 82.25 | 79.00 | 73.25 | ||||||
| no | 81.08 | 80.94 | 80.64 | 80.14 | 76.94 | 71.56 | ||||||
| pl | 84.00 | 83.97 | 84.06 | 83.39 | 81.28 | 76.83 | ||||||
| pt | 82.42 | 82.50 | 82.00 | 81.58 | 79.36 | 72.97 | ||||||
| ro | 89.36 | 89.39 | 89.22 | 89.14 | 87.03 | 82.06 | ||||||
| ru | 88.97 | 89.17 | 89.08 | 88.69 | 87.00 | 82.17 | ||||||
| sv | 78.06 | 77.86 | 78.00 | 77.47 | 74.83 | 69.53 | ||||||
| th | 81.61 | 81.64 | 81.14 | 80.36 | 76.97 | 68.17 | ||||||
| tr | 81.31 | 81.44 | 81.22 | 80.69 | 77.39 | 71.33 | ||||||
| uk | 88.56 | 88.42 | 88.61 | 87.89 | 85.86 | 80.72 | ||||||
| vi | 86.64 | 86.72 | 86.69 | 85.92 | 83.17 | 77.86 | ||||||
| zh | 78.97 | 79.06 | 78.86 | 78.08 | 74.81 | 67.36 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| average | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
| ar | 76.25 | 76.17 | 76.17 | 75.19 | 73.61 | 68.25 | ||||||
| bn | 69.00 | 69.00 | 68.94 | 68.75 | 66.19 | 59.39 | ||||||
| da | 88.53 | 88.39 | 88.25 | 87.58 | 85.69 | 80.25 | ||||||
| de | 92.47 | 92.56 | 92.42 | 92.31 | 90.64 | 87.19 | ||||||
| el | 73.33 | 73.50 | 73.47 | 73.39 | 72.75 | 68.61 | ||||||
| en | 78.58 | 78.61 | 78.36 | 77.86 | 76.44 | 72.58 | ||||||
| es | 86.28 | 86.39 | 86.33 | 85.78 | 84.03 | 80.17 | ||||||
| fi | 84.19 | 84.25 | 84.00 | 83.58 | 82.08 | 77.81 | ||||||
| fr | 90.89 | 90.86 | 90.75 | 90.14 | 88.61 | 84.89 | ||||||
| hi | 61.64 | 61.64 | 61.44 | 61.44 | 58.86 | 52.81 | ||||||
| id | 86.31 | 86.25 | 86.19 | 86.14 | 83.92 | 80.19 | ||||||
| it | 90.17 | 90.08 | 90.17 | 89.83 | 87.78 | 83.75 | ||||||
| ja | 88.50 | 88.64 | 88.64 | 87.92 | 86.22 | 81.64 | ||||||
| ko | 81.42 | 81.42 | 81.31 | 80.53 | 78.31 | 73.03 | ||||||
| nl | 82.47 | 82.44 | 82.36 | 81.94 | 79.97 | 76.00 | ||||||
| no | 83.75 | 83.81 | 83.78 | 83.28 | 80.39 | 75.39 | ||||||
| pl | 84.61 | 84.67 | 84.47 | 83.81 | 82.36 | 78.56 | ||||||
| pt | 83.94 | 83.83 | 83.64 | 83.25 | 81.72 | 77.00 | ||||||
| ro | 91.31 | 91.14 | 91.22 | 91.03 | 89.44 | 85.67 | ||||||
| ru | 90.64 | 90.58 | 90.31 | 90.19 | 88.50 | 84.86 | ||||||
| sv | 78.28 | 78.22 | 78.06 | 77.50 | 75.97 | 71.42 | ||||||
| th | 81.94 | 81.89 | 81.94 | 81.28 | 79.25 | 73.61 | ||||||
| tr | 82.67 | 82.72 | 82.39 | 81.92 | 80.36 | 76.06 | ||||||
| uk | 89.53 | 89.53 | 89.31 | 88.81 | 88.00 | 84.69 | ||||||
| vi | 88.06 | 87.94 | 87.61 | 87.47 | 85.92 | 82.00 | ||||||
| zh | 79.22 | 79.06 | 78.83 | 78.50 | 76.11 | 70.22 | ||||||
| Language - Dimension |
| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||
| average | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
| de | 85.70 | 85.40 | 84.90 | 84.70 | 81.10 | 79.30 | ||||||
| en | 89.40 | 89.60 | 88.70 | 88.90 | 86.50 | 83.00 | ||||||
| es | 85.90 | 85.90 | 86.00 | 85.70 | 84.00 | 80.80 | ||||||
| fr | 85.10 | 84.70 | 84.80 | 85.30 | 83.30 | 79.10 | ||||||
| it | 85.80 | 85.50 | 86.10 | 85.80 | 83.70 | 81.10 | ||||||
| ko | 82.10 | 81.90 | 82.10 | 80.10 | 77.90 | 71.90 | ||||||
| pl | 86.50 | 86.70 | 86.40 | 86.30 | 84.80 | 80.40 | ||||||
| ru | 81.10 | 81.20 | 80.80 | 81.40 | 77.80 | 74.00 | ||||||
| tr | 83.70 | 84.10 | 83.50 | 83.10 | 80.00 | 75.80 | ||||||
| zh | 83.40 | 83.50 | 82.70 | 81.90 | 78.90 | 73.10 | ||||||
| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||
| average | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
| de | 86.20 | 86.40 | 86.10 | 85.90 | 85.20 | 81.50 | ||||||
| en | 90.50 | 90.50 | 91.10 | 90.40 | 90.10 | 86.80 | ||||||
| es | 87.00 | 86.80 | 86.70 | 86.90 | 86.10 | 81.80 | ||||||
| fr | 85.20 | 84.90 | 85.20 | 85.00 | 84.20 | 81.70 | ||||||
| it | 88.00 | 88.00 | 87.70 | 87.60 | 85.70 | 83.20 | ||||||
| ko | 82.10 | 82.20 | 82.40 | 82.80 | 80.30 | 77.50 | ||||||
| pl | 88.50 | 88.70 | 88.60 | 88.30 | 85.90 | 83.60 | ||||||
| ru | 83.10 | 83.00 | 82.70 | 82.80 | 80.80 | 78.10 | ||||||
| tr | 85.60 | 85.60 | 85.90 | 85.90 | 84.50 | 79.80 | ||||||
| zh | 84.10 | 84.10 | 83.80 | 82.80 | 80.90 | 76.20 | ||||||
| Dataset - Dimensions |
| Retrieval - nDCG@10 | ||||||||||||
| Average Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
| FiQA2018 | 41.93 | 42.04 | 41.79 | 40.96 | 38.75 | 34.27 | ||||||
| NFCorpus | 32.89 | 33.01 | 32.72 | 32.32 | 30.20 | 25.69 | ||||||
| SciFact | 65.34 | 65.12 | 65.26 | 64.57 | 62.51 | 57.17 | ||||||
| SCIDOCS | 18.90 | 18.85 | 18.84 | 18.46 | 16.98 | 13.84 | ||||||
| CQADupstackRetrieval | 41.20 | 41.17 | 41.06 | 40.47 | 37.97 | 32.42 | ||||||
| Touche2020 | 23.94 | 24.08 | 24.23 | 24.57 | 25.82 | 23.26 | ||||||
| TRECCOVID | 76.73 | 76.54 | 75.98 | 76.01 | 73.10 | 67.47 | ||||||
| FEVER | 84.96 | 84.99 | 84.83 | 84.37 | 81.40 | 69.26 | ||||||
| HotpotQA | 60.14 | 60.13 | 59.71 | 58.06 | 52.14 | 38.97 | ||||||
| DBPedia | 37.44 | 37.42 | 36.98 | 36.12 | 33.21 | 26.34 | ||||||
| NQ | 57.17 | 57.19 | 57.02 | 56.30 | 53.49 | 45.42 | ||||||
| ClimateFEVER | 30.12 | 30.15 | 30.35 | 30.00 | 26.08 | 20.89 | ||||||
| MSMARCO | 37.47 | 37.43 | 37.45 | 37.16 | 35.74 | 32.47 | ||||||
| ArguAna | 43.55 | 43.50 | 43.53 | 42.83 | 41.06 | 37.21 | ||||||
| QuoraRetrieval | 88.14 | 88.13 | 88.06 | 87.88 | 87.10 | 85.19 | ||||||
| STS - Spearman correlation based on cosine similarity | ||||||||||||
| Average STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
| STS12 | 76.72 | 76.72 | 76.85 | 76.97 | 77.13 | 76.44 | ||||||
| STS13 | 79.90 | 79.89 | 79.98 | 80.48 | 80.00 | 78.69 | ||||||
| STS14 | 77.49 | 77.51 | 77.57 | 77.67 | 77.10 | 75.50 | ||||||
| STS15 | 86.42 | 86.43 | 86.46 | 86.58 | 86.21 | 84.58 | ||||||
| STS16 | 85.18 | 85.17 | 85.15 | 85.10 | 84.52 | 84.03 | ||||||
| STS17 | 87.87 | 87.89 | 87.74 | 87.31 | 86.91 | 85.36 | ||||||
| STS22 | 67.07 | 67.04 | 67.01 | 66.95 | 66.66 | 66.71 | ||||||
| BIOSSES | 83.00 | 83.03 | 82.65 | 82.33 | 81.18 | 78.27 | ||||||
| STSBenchmark | 86.87 | 86.68 | 86.86 | 86.72 | 86.26 | 85.14 | ||||||
| SICK-R | 82.38 | 82.38 | 82.36 | 82.27 | 81.87 | 80.89 | ||||||

### v2

@@
 Abstract
-Contrastive Language-Image Pretraining (CLIP) is a highly effective method for aligning images and texts in a shared embedding space.
-These models are widely used for tasks such as cross-modal information retrieval and multi-modal understanding.
-However, CLIP models often struggle with text-only tasks, underperforming compared to specialized text models.
-This performance disparity forces retrieval systems to rely on separate models for text-only and multi-modal tasks.
-In this work, we build upon our previous model, jina-clip-v1, by introducing a refined framework that utilizes multi-task, multi-stage contrastive learning across multiple languages, coupled with an improved training recipe to enhance text-only retrieval.
-The resulting model, jina-clip-v2, outperforms its predecessor on text-only and multimodal tasks, while adding multilingual support, better understanding of complex visual documents and efficiency gains thanks to Matryoshka Representation Learning and vector truncation.
-The model performs comparably to the state-of-the-art in both multilingual-multimodal and multilingual text retrieval benchmarks, addressing the challenge of unifying text-only and multi-modal retrieval systems.
+Contrastive Language-Image Pretraining (CLIP) has been widely used for crossmodal information retrieval and multimodal understanding tasks.
+However, CLIP models are mainly optimized for crossmodal vision-language tasks and underperform in single-mode text tasks.
+Moreover, these models are often trained on English datasets and therefore lack multilingual understanding.
+Additionally, from a visual understanding perspective, previous CLIP-based models exhibit insufficient understanding of visually rich documents.
+In this work, we propose jina-clip-v2, a contrastive vision-language model trained on text pairs, triplets and image-text pairs via a multi-task and multi-stage contrastive learning paradigm in order to support both text-only and crossmodal tasks.
+We employ a multilingual text encoder and expand the training dataset to include multilingual texts from 29 non-English languages, including Hindi, Chinese, German, French, and others, as well as images of visually-rich documents.
+We evaluate the model’s performance and show that jina-clip-v2 achieves notable improvements over state-of-the-art CLIP-based models in zero-shot text-only retrieval, semantic textual similarity, and crossmodal retrieval tasks in both English and multilingual settings.
+jina-clip-v2 also provides for flexibility in embedding dimensionality, enabling users to select the granularity of the representations.
+jina-clip-v2 is publicly available at https://huggingface.co/jinaai/jina-clip-v2.
 1 Introduction
-Contrastive text-image pre-training has become a widely adopted approach to building robust text-image alignment models.
-These models are particularly effective for tasks like cross-modal retrieval and zero-shot classification (Radford et al., 2021), demonstrating strong generalization capabilities.
-Conventional text-image pre-training excels at aligning text and image embeddings but struggles with aligning text-to-text embeddings (Koukounas et al., 2024).
-This is because cross-modal retrieval training is inadequate for text-only retrieval.
-The texts available are typically image captions, which are generally short and information-poor, and training procedures don’t use standard techniques for training text embeddings, like hard negatives.
-(Zhang et al., 2024a)
+Contrastive text-image pretraining is a well-known architecture and training framework for building robust text-image alignment models (Radford et al., 2021; Zhai et al., 2023; Sun et al., 2023).
+These models are particularly effective for tasks like crossmodal retrieval and zero-shot classification, demonstrating strong generalization capabilities.
+However, although CLIP models produce general-purpose embeddings for both texts and images, they are woefully inadequate as a source of text embeddings (Koukounas et al., 2024).
+We believe this is due to the training regimen of CLIP models, which are typically trained with short and information-poor image captions (Zhang et al., 2024a) and omit standard techniques for finetuning meaningful text embeddings, e.g.
+training using hard negatives (Gillick et al., 2019; Qu et al., 2021).
 To overcome these limitations, Koukounas et al.
 (2024) introduced a multi-task, multi-stage contrastive learning approach to simultaneously align text-text and text-image embeddings.
-Their method involves three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2) and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3).
-The resulting model, jina-clip-v1, achieves strong performance on both the cross-modal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and text embedding MTEB Benchmark (Muennighoff et al., 2023).
-Despite its strengths, this model has several limitations.
-First, it is an English-only multimodal embedding model, which makes it unsuitable for multilingual document retrieval.
-Second, jina-clip-v1 struggles with images that contain text, tables, graphs, diagrams, and other complex content (Faysse et al., 2024).
-To address these challenges, we propose an enhanced text-image pre-training scheme.
-Our approach incorporates multilingual text-image and text-text pairs, along with text-image pairs containing text and other difficult-to-process complex visual structures.
-Additionally, we gradually increase image resolution during pre-training to balance computational efficiency and model performance.
-The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on cross-modal retrieval benchmarks against state-of-the-art models like NLLB-CLIP-SigLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the Retrieval and STS tasks of the multilingual MTEB benchmark (Muennighoff et al., 2023).
-Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on benchmarks for visually rich document retrieval, compared to jina-clip-v1.
-To reduce vector storage costs, we employ Matryoshka Representation Learning (Kusupati et al., 2024) to enable dimension truncation on the output vectors with minimal performance degradation.
+This method involved three stages: optimizing text-text and text-image embeddings for short image-caption and text-text pairs (stage 1), refining alignment using long text-text pairs and detailed image-caption pairs (stage 2), and further enhancing performance with text triplets containing hard negatives and detailed image-caption pairs (stage 3).
+The resulting model, jina-clip-v1, achieves strong performance on both the crossmodal CLIP Benchmark111https://github.com/LAION-AI/CLIP_benchmark and the text embedding MTEB Benchmark (Muennighoff et al., 2023).
+Despite its strengths, jina-clip-v1 (Koukounas et al., 2024) has several limitations.
+First, it is an English-only model, which makes it unsuitable for multilingual document retrieval.
+Second, jina-clip-v1 struggles with visually-rich images, e.g.
+those containing text, tables, graphs, and diagrams (Faysse et al., 2024).
+Moreover, jina-clip-v1 only supports fixed length embeddings, which can be wasteful when embedding vectors are much larger than required for a task (Kusupati et al., 2024).
+To address these challenges, we propose several enhancements to the pretraining scheme.
+For multilingual support, our approach uses a multilingual language model (Conneau et al., 2020) to initialize the text encoder, and incorporates multilingual text-image and text-text pairs during training.
+To improve performance on visually complex documents, we gradually increase image resolution during pretraining and use multimodal training pairs with complex visual structures.
+Finally, to enable users to select the granularity of their representations, we employ Matryoshka Representation Learning (Kusupati et al., 2024) and enable dimensional truncation on the output vectors with minimal performance degradation.
+The resulting model, jina-clip-v2 depicted in Figure 1, not only achieves competitive performance on crossmodal retrieval benchmarks against multilingual state-of-the-art models like NLLB-CLIP (Visheratin, 2023b) but also performs comparably to dedicated multilingual text embedding models like jina-embeddings-v3 (Sturua et al., 2024) on the retrieval and semantic textual similarity (STS) tasks of the MTEB benchmark (Muennighoff et al., 2023).
+Moreover, due to the inclusion of visually rich training data and the progressive increase in image resolution, jina-clip-v2 demonstrates significantly improved performance on ViDoRe, a benchmark for visually rich document retrieval (Faysse et al., 2024), compared to jina-clip-v1.
+In summary, our contributions are the following:
+1.
+Support for multiple languages.
+jina-clip-v2 achieves up to 67% higher scores on multilingual crossmodal performance, up to 60% on multilingual Retrieval performance and up to 43% on multilingual STS, compared jina-clip-v1.
+2.
+Improved performance on visual document retrieval. jina-clip-v2’s scores are 35% higher on the ViDoRe benchmark (Faysse et al., 2024) compared to jina-clip-v1.
+3.
+Flexible embedding dimensionality. Truncating jina-clip-v2 embeddings by up to 75% (from 1,024 dimensions down to 256) reduces performance by <1%.
 2 Related Work
-Radford et al.
-(2021) pioneered the dual encoder architecture trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pre-trained vision-language models.
-Building upon this foundation, Zhai et al.
-(2023) propose an alternative and more efficient sigmoid objective, while Sun et al.
-(2023) optimize the model’s dimensions and scale up both dataset and model sizes, achieving state-of-the-art performance on cross-modal tasks.
+Previous work (Jia et al., 2021; Li et al., 2021; Radford et al., 2021; Zhang et al., 2022) pioneered dual-encoder architectures trained contrastively on image-text pairs, demonstrating impressive zero-shot performance and laying the groundwork for pretrained vision-language models.
+Building on this foundation, Zhai et al.
+(2023) proposed an alternative and more efficient sigmoid objective, while Sun et al.
+(2023) optimized the model’s dimensions and scaled up both dataset and model sizes, achieving state-of-the-art performance on crossmodal tasks.
 Koukounas et al.
-(2024) enhance a CLIP-based retriever with strong text-to-text and cross-modal retrieval capabilities, though fall short in addressing retrieval scenarios that involve candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022).
+(2024) enhanced a CLIP-based retriever with strong text-to-text and crossmodal retrieval capabilities, however, their work fails to address retrieval scenarios involving candidate pools with heterogeneous modalities, due to the modality gap (Liang et al., 2022; Eslami & de Melo, 2024) and the modality bias (Lin et al., 2024).
 To mitigate the modality bias, Lin et al.
-(2024) take a different approach by fine-tuning a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks.
-Extensive research has also focused on extending CLIP’s capabilities to multilingual contexts.
+(2024) finetuned a multimodal LLM with modality-aware hard-negative mining, creating a universal multimodal retriever and maintaining competitive text-to-text retrieval performance across diverse tasks.
+In this work, we build on the training strategy proposed by Koukounas et al.
+(2024) and train our model using a text-to-text as well as a text-to-image contrastive InfoNCE (Van den Oord et al., 2018) loss.
+Extensive research has also focused on extending CLIP to languages beyond English.
 To address the lack of image-caption pairs in languages other than English, Carlsson et al.
-(2022) apply knowledge distillation to retrain the text encoder using machine translation data.
-NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
-The importance of good information retrieval in Retrieval-Augmented Generation (RAG) systems has motivated rapid progress in multilingual text embedding models.
-These models are typically based on either a robust encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020) or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023).
-Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, leveraging extensive multilingual training with instruction tuning and multi-task learning techniques, respectively.
-mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the context length.
+(2022) applied knowledge distillation to retrain the text encoder using machine-translated data.
+NLLB-CLIP (Visheratin, 2023a) leverages Locked-image Tuning (LiT) (Zhai et al., 2022) alongside the multilingual NLLB (Costa-jussà et al., 2022) text encoder, achieving state-of-the-art results in both retrieval and classification tasks.
+Multilingual text embedding models are typically based on either an encoder-only architecture, such as XLM-RoBERTa (Conneau et al., 2020), or a decoder-only multilingual large language model, such as Mistral 7B (Jiang et al., 2023).
+Multilingual E5 (Wang et al., 2024) and BGE-M3 (Chen et al., 2024) both use XLM-RoBERTa as their backbone, the first leveraging extensive multilingual training with instruction tuning, and the second employing multi-task learning techniques.
+mGTE (Zhang et al., 2024b) achieves comparable performance to BGE-M3 using a smaller transformer architecture, enhanced by RoPE (Su et al., 2023) to extend the input context length.
 Similarly, Sturua et al.
 (2024) employ RoPE to increase the model’s context length and apply LoRA tuning (Hu et al., 2021) to optimize embeddings for downstream tasks.
-3 Training
+Finally, Kusupati et al.
+(2024) propose a training technique, which they call Matryoshka Representation Learning, that enables embedding models to learn coarse-to-fine representations.
+These representations can be truncated during inference to match the requirements of downstream tasks, cutting down on potentially redundant vector dimensions.
+3 jina-clip-v2
+The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024).
+The text encoder is initialized with pretrained Jina-XLM-RoBERTa model weights.
+Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the multilingual XLM-RoBERTa (Conneau et al., 2020) model to a modern encoder-only architecture with Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and support for low-rank adaptation (Hu et al., 2021).
+Like jina-clip-v1, the image encoder is a pretrained EVA02 model (Fang et al., 2023b).
+We selected the L/14 pretrained model variant, which is similar in number of parameters to the text encoder.
+This model implementation includes 2D rotary positional embeddings and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022).
+Table 1 presents architectural details for both encoders.
 | Feature | Text Encoder | Image Encoder |
@@
 | Parameters | 561M | 304M |
-| Input Specification | 8,192 tokens (max) | 512×512 pixels |
+| Input Specification | 8,192 tokens (max) | resolution |
 | Output Dimensions | 1,024 | 1,024 |
@@
 | Additional Features | 89 languages supported | Patch size 14x14 |
-The jina-clip-v2 model uses the dual encoder architecture, introduced in the original CLIP (Radford et al., 2021) model and reused in jina-clip-v1 (Koukounas et al., 2024).
-One encoder is responsible for encoding texts in a 1024-dimensional embedding space, while a second encoder embeds images in the same embedding space.
-The text encoder is initialized with the pre-trained Jina-XLM-RoBERTa model weights.
-Introduced in jina-embeddings-v3 (Sturua et al., 2024), the Jina-XLM-RoBERTa model is a port of the XLM-RoBERTa (Conneau et al., 2020) weights to a modern encoder-only architecture that adds Flash Attention (Dao et al., 2022), rotary positional embeddings (Su et al., 2023) and LoRA (Hu et al., 2021).
-For the image encoder, we have opted for the EVA02 (Fang et al., 2023b) family of ViT models, similarly to jina-clip-v1.
-We have selected a relatively large pre-trained model, similar in size to the text encoder.
-This model implementation includes 2D rotary positional embeddings (Su et al., 2023) and a memory-efficient attention implementation based on xFormers (Lefaudeux et al., 2022).
-Architectural details for both encoders are presented in Table 1.
-3.1 Datasets
-Similarly to jina-clip-v1 (Koukounas et al., 2024), we construct four different training datasets.
-Dataset denotes a dataset of text pairs, denotes a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset.
-All datasets are multi-lingual.
-The text corpora and are introduced in Sturua et al.
-(2024) and contain data in 30 languages.
-The text pair corpus consists of query and target pairs from a diverse collection of text-pair datasets.
-The triplet text corpus is a high-quality dataset with hard negatives from various sources.
-Each training item contains one annotated positive and seven negative items.
-Regarding , we focus on Retrieval and Semantic Textual Similarity (STS) datasets for the scope of this work.
-Our multi-modal training data comes from multiple sources.
-For the English part, we rely on the DFN dataset (Fang et al., 2023a), a collection of 400M image-caption pairs, filtered using data filtering networks.
-For the rest of the languages, we collect 400M samples from CommonPool (Gadre et al., 2023), using multilingual CLIP to identify and discard low-quality items.
-All images are resized to .
+3.1 Training Data
+Similarly to jina-clip-v1 (Koukounas et al., 2024), we constructed four training datasets used in different stages of the training process.
+Dataset denotes a dataset of text pairs, a dataset of text samples with hard negatives, a multimodal short-caption dataset and a multimodal long-caption dataset.
+Both and contain data in 30 languages drawn from various existing datasets for text information retrieval and semantic similarity, and were introduced as training data for jina-embeddings-v3 (Sturua et al., 2024).
+consists of pairs of texts drawn from a diverse collection of datasets, and is a high-quality dataset with hard negatives from various sources, where each sample contains one annotated positive and seven negative items.
+Sturua et al.
+(2024) provide detailed information about these datasets and how they were curated.
+Our multimodal training datasets draw on multiple sources.
+We randomly sampled 400M image-caption pairs from the 2B in the DFN dataset (Fang et al., 2023a) to obtain a collections of image-text pairs in English.
+To this, we added non-English image-text pairs sampled from CommonPool (Gadre et al., 2023).
+We filtered this data by language, aspect ratio, and by removing any caption with less than five words.
+We used multilingual SigLIP (Zhai et al., 2023) to rank image-text pairs based on cosine similarity and kept the top scoring 50%.
+The result was an additional 400M image-text pairs.
+All images were resized to .
 These two large-scale datasets are included in and a small part with captions longer than 256 tokens is held out for .
-Inspired by visual QA applications (Faysse et al., 2024), We have diversified our training data with PDFs, scientific graphs, infographics and Wikipedia images.
-The following datasets are included in both and : DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024).
-For the QA datasets, to obtain a caption for an image, we concatenate query and answer for each sample.
-For the WIT dataset, we use the reference caption in and a concatenation of caption, page title, section tile, page description and section description for .
+In order to overcome the limitations of visual document understanding, we diversify our training data with PDFs, scientific graphs, infographics and Wikipedia images.
+Specifically, the short caption image-text dataset and the long-caption include the following training datasets: DocVQA (Mathew et al., 2021b), TatDQA (Zhu et al., 2022), InfographicsVQA (Mathew et al., 2021a), SciGraphQA (Li & Tajbakhsh, 2023), ArXivQA and ArXivCAP (Li et al., 2024), WIT (Srinivasan et al., 2021) and ViDoRe synthetic training data (Faysse et al., 2024).
+For the QA datasets, we concatenated the query and answer for each sample to construct a matching text in lieu of a caption.
+For the WIT dataset, we use the reference caption as corresponding text in and a concatenation of reference caption, page title, section title, page description and section description as matching text for .
 Finally, includes multilingual synthetic long captions, similar to how Koukounas et al.
@@
 3.2 Training Stages
-Following the strategy proposed in Koukounas et al.
-(2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
-Stage 1 focuses on aligning the multi-modal representations while also improving the text representations of the text encoder.
-We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes.
+Inspired by Koukounas et al. (2024), we employ a multi-task, multi-stage training approach, in order to optimize the model for two tasks simultaneously: text-image matching and text-text matching.
+Stage 1 focuses on aligning the multimodal representations while also improving the text representations of the text encoder.
+We train on and , with a small context length of 77 and an image resolution of to enable large batch sizes of .
+This is the most computationally expensive stage in our training pipeline, requiring approximately 10 days on 8 NVIDIA H100 GPUs to converge.
 Towards the end, when performance peaks, we switch to a higher image resolution of using positional embedding interpolation as a warm-up for the next stage.
@@
 We do one more positional embedding interpolation to increase the image resolution to and train on this resolution for the duration of stage 3.
-Appendix table 6 specifies the training details for each step.
 3.3 Loss Functions
@@
 Similarly to , this loss function is bidirectional.
-Equation 3 formulates the loss for each of the three stages.
+Equation 3 defines the loss for each of the three stages.
 In each stage, we optimize a joint loss function, i.e.
-the sum of two loss functions, one operating on text representations and one on multi-modal representations.
+the sum of two loss functions, one operating on text representations and one on multimodal representations.
+In equation 3, denotes a batch of samples drawn from the dataset , where .
 All stages optimize except Stage 3, which uses hard negatives in the text loss branch and thus calculates .
-In Equation 3, denotes a batch of text pairs drawn from the dataset of text pairs , while denotes a batch of text samples with negatives drawn from the hard negatives dataset .
-Similarly denotes a batch of text-image pairs drawn from the multimodal short-caption dataset , while denotes a batch of text-image pairs sampled from the multimodal long-caption dataset .
 3.4 Matryoshka Representation Learning
-In all stages of training, each loss function is evaluated on the full 1024-dimension output vectors of the two encoders as well as on truncated vectors of 64, 128, 256, 512 and 768 dimensions.
-In other words, each loss component on equation 3 is the sum of the same loss function evaluated at different embedding dimensionalities.
-This training technique, called Matryoshka Representation Learning, (Kusupati et al., 2024) trains the model to learn the most important features in the vector sequence.
-This makes it possible, at inference time, to truncate embedding vectors, thereby reducing storage and computation costs when comparing vectors, at a minor performance penalty.
+In every training stage, the loss component is recalculated at different dimensionalities: Loss is computed for the full 1024-dim output vector, and for truncated subsets of 64, 128, 256, 512 and 768 dimensions.
+Model weight adjustment proceeds on the basis of all those losses combined.
+This training technique, called Matryoshka Representation Learning (Kusupati et al., 2024), induces the model to encode its embeddings by increasing granularity.
+This makes it possible, at inference time, to truncate embedding vectors with a minor performance penalty.
 4 Evaluation
-We have evaluated the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual cross-modal retrieval benchmarks in Section 4.1, common text embedding tasks from the MTEB benchmark suite in Section 4.2, visual document retrieval tasks in Section 4.3, and an ablation study of jina-clip-v2’s matryoshka representations in Section 4.4.
-Detailed evaluation results are presented in the appendices.
-Detailed CLIP Benchmark results are given in Appendix tables 7, 8 and 9, MTEB results are in in Appendix 10.
-ViDoRe benchmark results are in Appendix 11.
-A full MRL evaluation can be found in tables 12, 13, 14, 15 in the appendices.
-4.1 Cross-modal Evaluation
-| Benchmark | CLIP Benchmark - EN | Crossmodal-3600 | Multilingual MS COCO | |||
-| Task Type | Zero-Shot Retrieval | |||||
-| Model - Metric | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
-| nllb-siglip-large | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
-| nllb-siglip-base | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
-| jina-clip-v1 | 77.75 | 87.65 | - | - | - | - |
-| jina-clip-v2stage 1 | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
-| jina-clip-v2stage 2 | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
-| jina-clip-v2 | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |
+We evaluate the performance of jina-clip-v2 on a selection of benchmarks: English and multilingual crossmodal retrieval benchmarks in Section 4.1, text embedding tasks from the MTEB benchmark suite in Section 4.2 and visual document retrieval tasks in Section 4.3.
+Finally, we present an ablation study of jina-clip-v2’s Matryoshka representations in Section 4.4.
+Detailed evaluation results are presented in Section A tables 7 to 26.
+4.1 Crossmodal Retrieval Evaluation
+| Benchmark | CLIP Benchmark | Crossmodal-3600 | XTD10 | ||||
+| Language | English | Multilingual | Multilingual | ||||
+| Task Type | Zero-Shot Retrieval | ||||||
+| Model | #Parameters | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 | T-I r@5 | I-T r@5 |
+| jina-clip-v1 | 223M | 77.75 | 87.65 | 16.93 | 19.82 | 31.01 | 36.89 |
+| nllb-siglip-base | 507M | 79.57 | 86.38 | 79.29 | 76.56 | 86.23 | 84.87 |
+| nllb-siglip-large | 1.2B | 81.54 | 88.15 | 82.07 | 80.16 | 87.60 | 85.37 |
+| jina-clip-v2stage 1 | 865M | 73.87 | 86.61 | 73.51 | 79.64 | 80.66 | 83.02 |
+| jina-clip-v2stage 2 | 865M | 79.81 | 89.57 | 84.13 | 84.12 | 86.11 | 86.45 |
+| jina-clip-v2 | 865M | 79.09 | 89.73 | 81.43 | 83.23 | 84.87 | 86.03 |
 T-I r@5: Text to Image Recall@5 [%] I-T r@5: Image to Text Recall@5 [%]
-We evaluate jina-clip-v2 on a set of English and multilingual cross-modal tasks, conducting a comparative analysis against both its predecessor jina-clip-v1 and the NLLB-CLIP-SigLIP (Visheratin, 2023b) large and base variants, a top-performing multilingual CLIP model.
-Additionally, we track the model’s performance at different training stages to evaluate the effectiveness of different parts of the training protocol.
-Measuring 865M parameters, our model sits between the two NLLB-CLIP-SigLIP versions in terms of size: nllb-siglip-base (507M parameters, 41% smaller than ours) and nllb-siglip-large (1.2B parameters, 39% larger than ours).
-For English zero-shot image-text and text-image retrieval, we conduct evaluations on Flickr30K (Young et al., 2014) and MS COCO Captions (Chen et al., 2015), both of which are included in the CLIP Benchmark.
-For multilingual cross-modal tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and multilingual MS COCO (Chen et al., 2015) datasets.
-As shown in Table 2, jina-clip-v2 demonstrates strong performance across both English and multilingual benchmarks while showing clear progression through its training stages.
-On the English CLIP Benchmark, the final model achieves 79.09% and 89.73% on text-to-image and image-to-text retrieval respectively, showing significant improvement from stage 1’s 73.87% and 86.61%.
-While nllb-siglip-large leads in text-to-image retrieval at 81.54%, jina-clip-v2 outperforms its predecessor jina-clip-v1 and surpasses both NLLB-CLIP-SigLIP variants in image-to-text retrieval by 2.1% and 1.6% respectively.
-On the multilingual front, we obtain competitive results on Crossmodal-3600, achieving 81.43% and 83.23% on text-to-image and image-to-text retrieval respectively, approaching the performance of nllb-siglip-large.
-On multilingual MS COCO (Chen et al., 2015), jina-clip-v2 performs comparably to NLLB-CLIP-SigLIP models, reaching 84.87% and 86.03% on text-to-image and image-to-text retrieval.
-The improvement from Stage 1 to the final model is notable across all benchmarks, with gains of up to 10.6 percentage points on Crossmodal-3600’s text-to-image retrieval, demonstrating the effectiveness of the training strategy.
+We evaluate jina-clip-v2 on a set of English and multilingual crossmodal tasks, conducting a comparative analysis with both jina-clip-v1 and the state-of-the-art NLLB-CLIP (Visheratin, 2023b) large and base SigLIP variants.
+Additionally, we track the model’s performance at all training stages to evaluate the effectiveness of the training protocol.
+For English zero-shot image-text and text-image retrieval, we conduct evaluations on the CLIP Benchmark, which includes Flickr30K (Young et al., 2014) and COCO Captions (Chen et al., 2015).
+For multilingual crossmodal retrieval tasks, we assess performance on the Crossmodal-3600 (Thapliyal et al., 2022) and XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) datasets.
+Table 2 demonstrates the strong performance of jina-clip-v2 across both English and multilingual benchmarks.
+On the English CLIP Benchmark, our model outperforms jina-clip-v1 and while nllb-siglip-large leads in text-to-image retrieval, jina-clip-v2 outperforms both NLLB-CLIP variants in image-to-text retrieval.
+On multilingual multimodal retrieval, we obtain competitive results, approaching the performance of nllb-siglip-large in text-to-image retrieval and surpassing it in image-to-text retrieval.
+Regarding stage progression, performance improves considerably between Stage 1 and Stage 2 on all benchmarks, but drops slightly from Stage 2 to Stage 3.
+We hypothesize that there is a trade-off between crossmodal and text-only retrieval performance, and during stage 3 training, where focus shifts towards text embedding training with hard negatives, the model falls slightly behind on crossmodal retrieval.
 4.2 Text Retrieval and STS Evaluation
-| Benchmark | Retrieval | STS | ||
-| Metric | nDCG@10 [%] | Spearman Correlation [%] | ||
-| Model - Language | EN | Multilingual | EN | Multilingual |
-| nllb-siglip-large | 24.91 | - | 74.89 | - |
-| jina-embeddings-v3 | 53.87 | 72.58 | 85.80 | 69.81 |
-| jina-clip-v1 | 48.33 | - | 80.92 | - |
-| jina-clip-v2stage 1 | 40.91 | - | 79.67 | - |
-| jina-clip-v2stage 2 | 43.17 | - | 80.33 | - |
-| jina-clip-v2 | 49.32 | 69.85 | 81.29 | 67.77 |
-Table 3 presents the evaluation on MTEB tasks, demonstrating the effectiveness of jina-clip-v2’s text encoder for both retrieval and semantic textual similarity (STS) tasks.
-On English retrieval tasks (nDCG@10), jina-clip-v2 achieves an average score of 49.32, a significant improvement over nllb-siglip-large’s 24.91 and a slight improvement over jina-clip-v1’s 48.33, though not reaching jina-embeddings-v3’s score of 53.87.
-For multilingual retrieval, the model achieves 69.85, approaching jina-embeddings-v3’s performance of 72.58.
-In STS tasks measured by Spearman Correlation, jina-clip-v2 demonstrates competitive performance with 81.29 for English tasks, surpassing nllb-siglip-large’s 74.89 and comparable to jina-clip-v1’s 80.92, though below jina-embeddings-v3’s 85.8.
-On multilingual STS tasks, the model achieves 67.77, slightly below jina-embeddings-v3’s 69.8.
+| Retrieval nDCG@10 [%] | |||||||||
+| Model | #Parameters∗ | en | zh | hi | de | fr | es | jp | ru |
+| jina-clip-v1 | 137M | 47.04 | 8.09 | 8.73 | 33.77 | 39.99 | 39.77 | 10.77 | 4.78 |
+| nllb-siglip-base | 414M | 24.69 | 27.04 | 39.18 | 22.67 | 28.30 | 30.06 | 30.73 | 32.68 |
+| nllb-siglip-large | 767M | 27.83 | 32.96 | 43.22 | 32.29 | 38.74 | 39.92 | 35.53 | 38.87 |
+| jina-embeddings-v3 | 572M | 48.34 | 64.27 | 65.18 | 63.59 | 64.09 | 62.97 | 68.47 | 66.46 |
+| jina-clip-v2stage 1 | 561M | 42.02 | 55.11 | 62.03 | 57.56 | 58.56 | 59.86 | 63.34 | 60.74 |
+| jina-clip-v2stage 2 | 561M | 43.47 | 57.66 | 62.19 | 59.16 | 61.05 | 61.51 | 65.13 | 63.66 |
+| jina-clip-v2 | 561M | 46.46 | 60.47 | 63.08 | 56.66 | 61.16 | 61.48 | 65.48 | 64.73 |
+| STS Spearman correlation based on cosine similarity | |||||||||
+| Model | #Parameters∗ | en | zh | hi | de | fr | es | jp | ru |
+| jina-clip-v1 | 137M | 81.35 | 22.97 | 36.09 | 50.04 | 67.97 | 70.05 | 55.83 | 39.77 |
+| nllb-siglip-base | 414M | 72.57 | 33.36 | 68.11 | 39.58 | 63.94 | 65.28 | 77.39 | 57.55 |
+| nllb-siglip-large | 767M | 74.73 | 39.15 | 73.78 | 53.45 | 72.38 | 69.68 | 79.25 | 62.84 |
+| jina-embeddings-v3 | 572M | 81.68 | 54.60 | 83.62 | 75.32 | 81.11 | 80.78 | 81.03 | 77.16 |
+| jina-clip-v2stage 1 | 561M | 80.01 | 52.74 | 83.35 | 74.84 | 79.06 | 80.42 | 81.48 | 75.62 |
+| jina-clip-v2stage 2 | 561M | 80.76 | 53.37 | 84.30 | 75.64 | 80.19 | 80.51 | 81.46 | 76.18 |
+| jina-clip-v2 | 561M | 81.58 | 55.08 | 79.55 | 76.27 | 81.70 | 81.89 | 82.01 | 77.79 |
+∗ Refers to the parameters of the text tower
+Table 3 presents the results of evaluations on the retrieval and semantic textual similarity (STS) tasks from MTEB (Muennighoff et al., 2023), both in English-only and multilingual contexts.
+jina-clip-v2 significantly improves on the performance of nllb-siglip-large in both English and multilingual text tasks, demonstrating the value of explicit text retrieval training.
+Compared to jina-clip-v1, performance is significantly better on multilingual tasks, and comparable on English-only tasks.
+Compared to the text-only frontier model jina-embeddings-v3, performance is lackluster on retrieval tasks.
+This suggests that multimodal training is hampering text-only performance.
 4.3 Visual Document Retrieval
-| Model | Average (nDCG@5) [%] |
-| jina-clip-v1 | 17.72 |
-| nllb-siglip-large | 46.55 |
-| jina-clip-v2 | 52.65 |
-| jina-clip-v2stage 1 | 37.37 |
-| jina-clip-v2stage 2 | 47.76 |
-Visual document retrieval requires vision models to capture fine-grained information from images to accurately match documents with queries.
-This task is particularly challenging as it deals with high-resolution document images, which differ significantly from typical image-caption datasets.
-As shown in table 4, jina-clip-v2 achieves state-of-the-art performance on the ViDoRe Benchmark (Faysse et al., 2024) for visual document understanding, with an average nDCG@5 score of 52.65%, significantly outperforming both its predecessor jina-clip-v1 (17.72%) and nllb-siglip-large (46.55%).
-The progression through training stages demonstrates substantial improvements, advancing from 37.37% in stage 1 to 47.76% in stage 2, before reaching peak performance in the final model.
-These results highlight the effectiveness of our training strategy and the impact of the new datasets introduced for document retrieval tasks.
+| Task |
+| Retrieval - nDCG@5 [%] | ||||||||||||||||||
+| ArxivQ | 64.92 | 58.78 | 66.14 | 25.40 | 30.44 | 43.2 | ||||||||||||
+| DocQ | 24.64 | 21.16 | 24.81 | 11.90 | 23.82 | 30.3 | ||||||||||||
+| InfoQ | 57.90 | 56.62 | 56.63 | 35.50 | 59.87 | 64.1 | ||||||||||||
+| TabF | 45.91 | 37.28 | 43.18 | 20.20 | 70.68 | 58.1 | ||||||||||||
+| TATQ | 30.25 | 12.09 | 25.79 | 3.30 | 20.00 | 26.2 | ||||||||||||
+| Shift | 34.07 | 28.41 | 31.54 | 3.80 | 30.79 | 18.7 | ||||||||||||
+| AI | 68.07 | 32.95 | 55.64 | 15.20 | 47.90 | 62.5 | ||||||||||||
+| Energy | 62.15 | 49.15 | 60.97 | 19.70 | 64.94 | 65.7 | ||||||||||||
+| Gov | 68.97 | 37.81 | 55.56 | 21.40 | 58.62 | 66.1 | ||||||||||||
+| Health | 69.05 | 39.50 | 57.37 | 20.80 | 58.43 | 79.1 | ||||||||||||
+| Average | 52.65 | 37.37 | 47.76 | 17.72 | 46.55 | 51.4 | ||||||||||||
+Visual document retrieval challenges vision-language models to capture fine-grained information from images, including embedded text, and accurately match these with text queries or documents.
+This task is particularly challenging as it deals with high-resolution document images with complex content, which differs significantly from processing typical image-caption datasets.
+As shown in Table 4, jina-clip-v2 outperforms other state-of-the-art CLIP models on the ViDoRe Benchmark for visual document understanding (Faysse et al., 2024), with an average nDCG@5 score of 52.65%, toping jina-clip-v1, nllb-siglip-large and siglip-so400m-patch14-384 (Zhai et al., 2023; Alabdulmohsin et al., 2024).
+The progression through training stages brings substantial improvements, highlighting the effectiveness of gradually increasing the image resolution.
+We further investigate how image resolution affects visual document retrieval in Section 5.1.
 4.4 Matryoshka Representation Learning
@@
 | Text-Image Retrieval - Recall@5 [%] | ||||||||||||
-| Avg CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
-| Avg Crossmodal-3600 | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
-| Avg Multilingual MS COCO | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
+| CLIP Benchmark | 79.10 | 79.12 | 78.93 | 78.32 | 75.90 | 70.51 | ||||||
+| Crossmodal-3600 (Thapliyal et al., 2022) | 81.43 | 82.35 | 82.31 | 81.75 | 78.17 | 72.52 | ||||||
+| XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) | 84.87 | 84.85 | 84.60 | 84.32 | 81.80 | 77.85 | ||||||
 | Image-Text Retrieval - Recall@5 [%] | ||||||||||||
-| Avg CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
-| Avg Crossmodal-3600 | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
-| Avg Multilingual MS COCO | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
+| CLIP Benchmark | 89.73 | 89.60 | 89.55 | 89.35 | 87.48 | 83.20 | ||||||
+| Crossmodal-3600 (Thapliyal et al., 2022) | 83.23 | 83.26 | 83.21 | 82.81 | 80.54 | 75.37 | ||||||
+| XTD10 (Aggarwal & Kale, 2020; Rajendran et al., 2016) | 86.03 | 86.02 | 86.02 | 85.84 | 84.37 | 81.02 | ||||||
 | Text-Text Retrieval - nDCG@10 [%] | ||||||||||||
-| Avg Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
+| EN Classic MTEB (Muennighoff et al., 2023) Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
 | Semantic Textual Similarity - Spearman Correlation [%] | ||||||||||||
-| Avg STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
-Table 5 presents the impact of embedding truncation on the performance of our model.
-The table is divided into 4 task categories: text-to-image retrieval, image-to-text retrieval, text-to-text retrieval and semantic textual similarity (STS).
-The results show that performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks.
-At 256 dimensions, representing a 75% reduction in embedding size, the model effectively preserves essential semantic information.
-Substantial performance degradation is only observed at very low dimensions (128 and 64), indicating the model’s very robust acquisition of image and text features.
+| EN Classic MTEB (Muennighoff et al., 2023) STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
+Table 5 presents the impact of embedding truncation on the crossmodal and text-only tasks.
+Performance remains highly stable when reducing dimensions from 1024 to 256, with minimal degradation (typically less than 1%) across all evaluation tasks.
+At 256 dimensions, representing a 75% reduction in embedding size, the model preserves effectively all essential semantic information.
+Substantial performance degradation is only observed at dimensions 128 and 64.
 5 Analysis
-5.1 The Role of Image Resolution in Complex Document Retrieval
-Our empirical study demonstrates that image resolution plays a critical role in retrieving visually rich documents.
-To better understand the relationship between image resolution and model performance on such documents, we have conducted a targeted experiment.
-We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of (224, 224).
-This model checkpoint was trained on 1.45 billion image-caption pairs and 1.45 billion text-text pairs.
+5.1 The Role of Image Resolution in Visual Document Retrieval
+Our analysis of performance at various stages of training demonstrates that image resolution plays a critical role in retrieving visually rich documents.
+To better understand the relationship between image resolution and model performance on such documents, we conduct a targeted experiment.
+We checkpointed jina-clip-v2 at the end of Stage 1, when it had only seen images with a resolution of .
 We conducted four runs, keeping the image resolution to for the first run and increasing to , , and for three additional runs.
-Each run is trained for 3500 steps using the same visually rich training set and hyperparameters.
-The models are then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents, based on text queries.
-We have plotted the results in Figure 2.
-Unsurprisingly, increasing image resolution has a positive impact on linking captions to visually rich documents.
+Each run was trained for 3500 steps using the same visually rich training set and hyperparameters.
+The models were then evaluated on the ViDoRe benchmark (Faysse et al., 2024), which comprises 10 datasets designed for retrieving visually rich documents with text queries.
+We plot the results in Figure 2.
+Unsurprisingly, increasing image resolution has a positive impact on linking queries to visually rich documents.
 The most significant improvement occurs when resolution increases from to , with the average nDCG@5 score across 10 benchmarks rising from to .
 A further increase in resolution to also brings a noticeable gain in performance.
-Considering that the model is trained with a patch size of 14, increasing the resolution beyond significantly raises the number of patches, a increase when moving from to .
-This leads to substantial additional computational overhead while yielding only marginal improvement, with nDCG@5 increasing by just .
-Therefore, we identify as an optimal image resolution, offering a good balance between performance and computational efficiency.
-5.2 Unified Batch vs Multi-Task Learning
-Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multi-modal alignment task for images and text, and second text-only task.
+Considering that the model is trained with a patch size of 14, increasing the resolution beyond raises the number of patches quadratically, for example, by a factor of when increasing the resolution to .
+This large increase in computing costs yields only marginal improvements, with nDCG@5 increasing by just .
+We consider the optimal image resolution for this model, with a reasonable balance between performance and cost.
+5.2 Unified Batch Training vs Multi-Task Learning
+Our training approach, as outlined in Section 3, involves two tasks at each stage, both optimized in a multi-task learning framework: a multimodal alignment task for images and text, and a second text-only task.
 Both use the InfoNCE (Van den Oord et al., 2018) objective.
 Regardless of whether the loss is computed on image-text or text-text pairs, the same loss function is applied to pairs of vectors.
-This raises the question: Would unifying the two batches of pairs into a single batch, irrespective of modality, and optimizing a single contrastive objective, offer any benefits?
-We refer to this approach as the unified batch technique, in contrast to the multi-task learning paradigm.
-We present the training paradigm in Figure 3.
-The motivation is two-fold: first, simplicity of computation and secondly, modality gap mitigation (Liang et al., 2022) by using in-batch negatives across modalities.
-Compared to multi-task training, the unified batch approach utilizes a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space.
-Additionally, we integrate a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representation.
-However, our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to cross-modal tasks.
+This raises the question: Is there any value in merging the image-text pair and text-text pair data into mixed modality batches and optimizing a single contrastive objective?
+We refer to this approach as the Unified Batch Training technique, in contrast to the Multi-Task Learning paradigm.
+The motivation is two-fold: simplicity of computation, and the possibility of mitigating the modality gap (Liang et al., 2022) by using in-batch negatives across modalities.
+Compared to multi-task training, the unified batch approach would use a single InfoNCE loss with a shared temperature value, attempting to force all modalities to align within the same embedding space.
+For this investigation, we also integrated a SimCLR-like self-supervised learning method (Chen et al., 2020) into the contrastive training, aiming to provide more robust image representations.
+A schematic representation of the unified batch technique is given in Figure 3 in Section A.
+Our Stage 1 experiments show that this technique does bring some early improvements, but fails beyond a certain point, at least with regard to crossmodal tasks.
 This was reflected in the slow decay of the loss temperature (our temperature is a trainable parameter in this case) and its eventual stabilization around .
-This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training, leading to better performance on cross-modal tasks.
+This is in contrast to the multi-task learning approach, where the temperature value decreases more rapidly and keeps decreasing throughout training down to , leading to better performance on crossmodal tasks.
 We hypothesize that this limitation arises from the fundamental information asymmetry between the visual and textual modalities (Schrodi et al., 2024).
-Images include fine-grained details of information whereas the corresponding textual descriptions include more coarse-grained level of information (Schrodi et al., 2024; Chen et al., 2023b).
+Images are very information-dense, full of small details, while the corresponding textual descriptions are far less detailed (Schrodi et al., 2024; Chen et al., 2023b).
 In temperature-scaled contrastive learning, the temperature parameter controls the hardness of the loss function with respect to the negative samples, i.e., the level of penalties on the hard negative samples (Wang & Liu, 2021).
 From an information-theory perspective, the level of penalty on the hard negative text-text samples theoretically differs from that of image-text samples due to the information gap problem.
-Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is potentially theoretically suboptimal.
+Consequently, as supported by our experiments, enforcing a unified temperature parameter across both modalities is empirically suboptimal.
 6 Conclusion
-In this work we present an enhanced strategy for training dual-encoder vision-language embedding models for cross-modal and text retrieval tasks, using contrastive learning.
-We introduce improvements to the pipeline, namely support for multiple languages, Matryoshka Representation Learning, and finer visual perception.
-We have trained and released a new model, jina-clip-v2, which has strong cross-modal and text retrieval performance on standard benchmarks.
-Finally, we have conducted analyses of three important considerations for CLIP models going forward: the ability to understand complex visual inputs performance how the image resolution affects it, and the effects of efforts to mitigate the modality gap.
+This work presents an enhanced strategy for training dual-encoder vision-language embedding models for multilingual crossmodal as well as text retrieval and semantic text similarity tasks using contrastive learning.
+We introduce improvements to the training, namely support for multiple languages, Matryoshka Representation Learning, and visually-rich document understanding.
+We have used these improvements to train and release a new model, jina-clip-v2, with strong crossmodal and text-only performance on standard benchmarks.
+Finally, we have presented in this paper analyses of two important considerations for CLIP models going forward: the effects of the modality gap and the role of the image resolution in understanding complex visual inputs.
 Appendix A Appendix
@@
 | Number of GPUs - H100s 80GB | 8 | 8 | 8 |
+| Input resolution | |||
+| Patch size | |||
 | LR schedule | cosine decay | ||
@@
 | Weight decay | 0.02 | ||
-| Input resolution | (224…384, 224…384) | (384, 384) | (512, 512) |
-| Patch size | (14, 14) | ||
 | Numerical precision | bfloat16 | ||
@@
 | Language - Model |
-| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
-| average | 81.43 | 73.51 | 84.13 | 82.07 | 79.29 | |||||||||
-| ar | 73.56 | 66.22 | 76.89 | 78.92 | 76.94 | |||||||||
-| bn | 63.78 | 50.19 | 68.11 | 75.19 | 74.19 | |||||||||
-| da | 85.39 | 76.53 | 87.97 | 87.14 | 86.64 | |||||||||
-| de | 91.25 | 84.64 | 92.42 | 89.56 | 87.25 | |||||||||
-| el | 75.03 | 68.53 | 77.92 | 77.83 | 71.97 | |||||||||
-| en | 75.83 | 69.78 | 77.78 | 73.11 | 72.22 | |||||||||
-| es | 83.64 | 78.28 | 86.28 | 82.64 | 79.97 | |||||||||
-| fi | 82.83 | 75.28 | 85.89 | 86.42 | 81.44 | |||||||||
-| fr | 88.78 | 83.50 | 90.67 | 87.86 | 85.58 | |||||||||
-| hi | 55.25 | 42.03 | 59.47 | 60.31 | 58.08 | |||||||||
-| id | 84.22 | 73.69 | 86.94 | 86.31 | 84.17 | |||||||||
-| it | 88.33 | 81.19 | 89.64 | 85.94 | 82.67 | |||||||||
-| ja | 87.03 | 77.28 | 90.03 | 86.06 | 83.14 | |||||||||
-| ko | 78.81 | 71.81 | 83.22 | 78.75 | 75.47 | |||||||||
-| nl | 82.56 | 75.72 | 84.47 | 81.69 | 78.86 | |||||||||
-| no | 81.08 | 71.97 | 83.08 | 82.69 | 79.97 | |||||||||
-| pl | 84.00 | 79.33 | 86.50 | 82.72 | 78.61 | |||||||||
-| pt | 82.42 | 76.17 | 85.19 | 82.69 | 79.44 | |||||||||
-| ro | 89.36 | 82.92 | 92.22 | 90.03 | 86.17 | |||||||||
-| ru | 88.97 | 82.97 | 91.11 | 86.44 | 83.78 | |||||||||
-| sv | 78.06 | 71.33 | 80.56 | 79.33 | 76.17 | |||||||||
-| th | 81.61 | 70.92 | 85.08 | 81.14 | 78.83 | |||||||||
-| tr | 81.31 | 75.31 | 84.53 | 83.47 | 81.00 | |||||||||
-| uk | 88.56 | 82.28 | 89.89 | 85.44 | 81.89 | |||||||||
-| vi | 86.64 | 76.31 | 89.06 | 85.56 | 82.56 | |||||||||
-| zh | 78.97 | 66.97 | 82.50 | 76.56 | 74.64 | |||||||||
-| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
-| average | 83.23 | 79.64 | 84.12 | 80.16 | 76.56 | |||||||||
-| ar | 76.25 | 72.14 | 76.67 | 75.86 | 73.69 | |||||||||
-| bn | 69.00 | 63.78 | 70.22 | 75.58 | 73.61 | |||||||||
-| da | 88.53 | 84.42 | 88.86 | 86.53 | 84.69 | |||||||||
-| de | 92.47 | 88.42 | 92.47 | 87.50 | 84.44 | |||||||||
-| el | 73.33 | 73.22 | 75.61 | 74.81 | 68.89 | |||||||||
-| en | 78.58 | 74.33 | 79.61 | 70.81 | 69.08 | |||||||||
-| es | 86.28 | 81.19 | 87.36 | 81.19 | 77.00 | |||||||||
-| fi | 84.19 | 81.17 | 85.64 | 84.25 | 78.75 | |||||||||
-| fr | 90.89 | 86.56 | 91.00 | 86.64 | 83.64 | |||||||||
-| hi | 61.64 | 54.39 | 61.56 | 61.89 | 59.83 | |||||||||
-| id | 86.31 | 83.64 | 87.64 | 84.33 | 81.89 | |||||||||
-| it | 90.17 | 84.31 | 90.53 | 83.50 | 78.64 | |||||||||
-| ja | 88.50 | 85.53 | 89.44 | 84.03 | 81.14 | |||||||||
-| ko | 81.42 | 79.22 | 83.00 | 76.75 | 74.17 | |||||||||
-| nl | 82.47 | 77.94 | 83.33 | 79.28 | 74.33 | |||||||||
-| no | 83.75 | 77.22 | 84.42 | 82.08 | 77.03 | |||||||||
-| pl | 84.61 | 83.72 | 85.69 | 79.58 | 75.39 | |||||||||
-| pt | 83.94 | 79.31 | 84.06 | 79.39 | 74.89 | |||||||||
-| ro | 91.31 | 87.39 | 92.08 | 88.83 | 83.58 | |||||||||
-| ru | 90.64 | 87.75 | 90.89 | 84.64 | 80.61 | |||||||||
-| sv | 78.28 | 77.53 | 80.11 | 76.58 | 72.31 | |||||||||
-| th | 81.94 | 80.22 | 84.11 | 79.56 | 76.28 | |||||||||
-| tr | 82.67 | 80.36 | 83.81 | 80.36 | 77.69 | |||||||||
-| uk | 89.53 | 86.42 | 89.97 | 83.22 | 78.67 | |||||||||
-| vi | 88.06 | 85.64 | 88.58 | 83.14 | 80.03 | |||||||||
-| zh | 79.22 | 74.94 | 80.42 | 73.83 | 70.22 | |||||||||
+| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||||
+| average | 81.43 | 73.51 | 84.13 | 16.93 | 82.07 | 79.29 | ||||||||||
+| ar | 73.56 | 66.22 | 76.89 | 0.19 | 78.92 | 76.94 | ||||||||||
+| bn | 63.78 | 50.19 | 68.11 | 0.11 | 75.19 | 74.19 | ||||||||||
+| da | 85.39 | 76.53 | 87.97 | 15.39 | 87.14 | 86.64 | ||||||||||
+| de | 91.25 | 84.64 | 92.42 | 37.42 | 89.56 | 87.25 | ||||||||||
+| el | 75.03 | 68.53 | 77.92 | 0.56 | 77.83 | 71.97 | ||||||||||
+| en | 75.83 | 69.78 | 77.78 | 76.17 | 73.11 | 72.22 | ||||||||||
+| es | 83.64 | 78.28 | 86.28 | 47.36 | 82.64 | 79.97 | ||||||||||
+| fi | 82.83 | 75.28 | 85.89 | 5.03 | 86.42 | 81.44 | ||||||||||
+| fr | 88.78 | 83.50 | 90.67 | 62.33 | 87.86 | 85.58 | ||||||||||
+| hi | 55.25 | 42.03 | 59.47 | 0.11 | 60.31 | 58.08 | ||||||||||
+| id | 84.22 | 73.69 | 86.94 | 13.14 | 86.31 | 84.17 | ||||||||||
+| it | 88.33 | 81.19 | 89.64 | 33.39 | 85.94 | 82.67 | ||||||||||
+| ja | 87.03 | 77.28 | 90.03 | 3.75 | 86.06 | 83.14 | ||||||||||
+| ko | 78.81 | 71.81 | 83.22 | 0.33 | 78.75 | 75.47 | ||||||||||
+| nl | 82.56 | 75.72 | 84.47 | 27.69 | 81.69 | 78.86 | ||||||||||
+| no | 81.08 | 71.97 | 83.08 | 15.61 | 82.69 | 79.97 | ||||||||||
+| pl | 84.00 | 79.33 | 86.50 | 7.39 | 82.72 | 78.61 | ||||||||||
+| pt | 82.42 | 76.17 | 85.19 | 31.97 | 82.69 | 79.44 | ||||||||||
+| ro | 89.36 | 82.92 | 92.22 | 17.89 | 90.03 | 86.17 | ||||||||||
+| ru | 88.97 | 82.97 | 91.11 | 2.19 | 86.44 | 83.78 | ||||||||||
+| sv | 78.06 | 71.33 | 80.56 | 15.22 | 79.33 | 76.17 | ||||||||||
+| th | 81.61 | 70.92 | 85.08 | 1.89 | 81.14 | 78.83 | ||||||||||
+| tr | 81.31 | 75.31 | 84.53 | 4.86 | 83.47 | 81.00 | ||||||||||
+| uk | 88.56 | 82.28 | 89.89 | 0.94 | 85.44 | 81.89 | ||||||||||
+| vi | 86.64 | 76.31 | 89.06 | 3.17 | 85.56 | 82.56 | ||||||||||
+| zh | 78.97 | 66.97 | 82.50 | 2.06 | 76.56 | 74.64 | ||||||||||
+| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||||
+| average | 83.23 | 79.64 | 84.12 | 19.82 | 80.16 | 76.56 | ||||||||||
+| ar | 76.25 | 72.14 | 76.67 | 0.31 | 75.86 | 73.69 | ||||||||||
+| bn | 69.00 | 63.78 | 70.22 | 0.11 | 75.58 | 73.61 | ||||||||||
+| da | 88.53 | 84.42 | 88.86 | 17.25 | 86.53 | 84.69 | ||||||||||
+| de | 92.47 | 88.42 | 92.47 | 39.75 | 87.50 | 84.44 | ||||||||||
+| el | 73.33 | 73.22 | 75.61 | 0.64 | 74.81 | 68.89 | ||||||||||
+| en | 78.58 | 74.33 | 79.61 | 79.06 | 70.81 | 69.08 | ||||||||||
+| es | 86.28 | 81.19 | 87.36 | 49.14 | 81.19 | 77.00 | ||||||||||
+| fi | 84.19 | 81.17 | 85.64 | 7.22 | 84.25 | 78.75 | ||||||||||
+| fr | 90.89 | 86.56 | 91.00 | 63.14 | 86.64 | 83.64 | ||||||||||
+| hi | 61.64 | 54.39 | 61.56 | 0.08 | 61.89 | 59.83 | ||||||||||
+| id | 86.31 | 83.64 | 87.64 | 16.75 | 84.33 | 81.89 | ||||||||||
+| it | 90.17 | 84.31 | 90.53 | 36.53 | 83.50 | 78.64 | ||||||||||
+| ja | 88.50 | 85.53 | 89.44 | 7.69 | 84.03 | 81.14 | ||||||||||
+| ko | 81.42 | 79.22 | 83.00 | 0.53 | 76.75 | 74.17 | ||||||||||
+| nl | 82.47 | 77.94 | 83.33 | 29.28 | 79.28 | 74.33 | ||||||||||
+| no | 83.75 | 77.22 | 84.42 | 18.53 | 82.08 | 77.03 | ||||||||||
+| pl | 84.61 | 83.72 | 85.69 | 10.00 | 79.58 | 75.39 | ||||||||||
+| pt | 83.94 | 79.31 | 84.06 | 33.33 | 79.39 | 74.89 | ||||||||||
+| ro | 91.31 | 87.39 | 92.08 | 18.83 | 88.83 | 83.58 | ||||||||||
+| ru | 90.64 | 87.75 | 90.89 | 3.33 | 84.64 | 80.61 | ||||||||||
+| sv | 78.28 | 77.53 | 80.11 | 16.50 | 76.58 | 72.31 | ||||||||||
+| th | 81.94 | 80.22 | 84.11 | 2.17 | 79.56 | 76.28 | ||||||||||
+| tr | 82.67 | 80.36 | 83.81 | 6.94 | 80.36 | 77.69 | ||||||||||
+| uk | 89.53 | 86.42 | 89.97 | 1.58 | 83.22 | 78.67 | ||||||||||
+| vi | 88.06 | 85.64 | 88.58 | 6.06 | 83.14 | 80.03 | ||||||||||
+| zh | 79.22 | 74.94 | 80.42 | 4.67 | 73.83 | 70.22 | ||||||||||
 | Language - Model |
-| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||
-| average | 84.87 | 80.66 | 86.11 | 87.60 | 86.23 | |||||||||
-| de | 85.70 | 80.40 | 86.80 | 88.30 | 87.00 | |||||||||
-| en | 89.40 | 84.00 | 89.40 | 89.40 | 88.80 | |||||||||
-| es | 85.90 | 82.70 | 87.40 | 88.20 | 86.70 | |||||||||
-| fr | 85.10 | 80.90 | 87.30 | 87.70 | 86.90 | |||||||||
-| it | 85.80 | 83.20 | 87.10 | 89.30 | 87.80 | |||||||||
-| ko | 82.10 | 78.30 | 83.00 | 85.20 | 83.60 | |||||||||
-| pl | 86.50 | 81.10 | 88.00 | 89.40 | 87.60 | |||||||||
-| ru | 81.10 | 79.30 | 82.10 | 83.40 | 82.40 | |||||||||
-| tr | 83.70 | 81.30 | 86.30 | 88.30 | 86.80 | |||||||||
-| zh | 83.40 | 75.40 | 83.70 | 86.80 | 84.70 | |||||||||
-| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||
-| average | 86.03 | 83.02 | 86.45 | 85.37 | 84.56 | |||||||||
-| de | 86.20 | 82.90 | 86.90 | 84.40 | 84.30 | |||||||||
-| en | 90.50 | 87.80 | 90.60 | 88.30 | 87.50 | |||||||||
-| es | 87.00 | 84.90 | 87.50 | 85.80 | 85.20 | |||||||||
-| fr | 85.20 | 82.20 | 85.00 | 86.30 | 84.40 | |||||||||
-| it | 88.00 | 83.10 | 87.70 | 86.70 | 85.60 | |||||||||
-| ko | 82.10 | 79.30 | 84.00 | 83.20 | 82.60 | |||||||||
-| pl | 88.50 | 84.50 | 88.90 | 86.80 | 86.60 | |||||||||
-| ru | 83.10 | 79.60 | 81.80 | 80.90 | 81.20 | |||||||||
-| tr | 85.60 | 83.20 | 85.90 | 87.10 | 85.40 | |||||||||
-| zh | 84.10 | 82.70 | 86.20 | 84.20 | 82.80 | |||||||||
-| Dataset - Model |
-| Retrieval - nDCG@10 | ||||||||||||||||
-| Average Retrieval | 49.33 | 53.87 | 40.91 | 43.17 | 24.92 | 48.33 | ||||||||||
-| ArguAna | 43.55 | 54.33 | 41.87 | 45.84 | 40.63 | 49.36 | ||||||||||
-| ClimateFEVER | 30.12 | 42.36 | 19.32 | 23.74 | 15.57 | 24.81 | ||||||||||
-| CQADupstackRetrieval | 41.20 | 42.36 | 41.57 | 42.09 | 21.06 | 40.92 | ||||||||||
-| DBPedia | 37.44 | 41.00 | 22.04 | 25.48 | 20.46 | 36.64 | ||||||||||
-| FEVER | 84.96 | 89.05 | 60.97 | 66.13 | 14.98 | 76.28 | ||||||||||
-| FiQA2018 | 41.93 | 47.35 | 41.87 | 40.97 | 12.90 | 38.27 | ||||||||||
-| HotpotQA | 60.14 | 64.67 | 49.28 | 47.00 | 22.38 | 61.89 | ||||||||||
-| MSMARCO | 37.47 | 40.82 | 23.06 | 27.05 | 13.35 | 36.91 | ||||||||||
-| NFCorpus | 32.89 | 36.62 | 30.95 | 32.11 | 18.97 | 33.52 | ||||||||||
-| NQ | 57.17 | 64.23 | 39.78 | 42.80 | 16.96 | 58.09 | ||||||||||
-| QuoraRetrieval | 88.14 | 89.09 | 87.12 | 86.92 | 80.71 | 87.88 | ||||||||||
-| SCIDOCS | 18.90 | 19.81 | 20.51 | 20.91 | 10.62 | 20.24 | ||||||||||
-| SciFact | 65.34 | 72.31 | 68.26 | 67.61 | 35.28 | 67.34 | ||||||||||
-| TRECCOVID | 76.73 | 77.72 | 51.29 | 58.63 | 38.15 | 71.61 | ||||||||||
-| Touche2020 | 23.94 | 26.30 | 15.72 | 20.21 | 11.75 | 21.15 | ||||||||||
-| STS - Spearman correlation based on cosine similarity | ||||||||||||||||
-| Average STS | 81.29 | 85.80 | 79.67 | 80.33 | 74.90 | 80.92 | ||||||||||
-| BIOSSES | 83.00 | 88.69 | 83.55 | 83.62 | 65.07 | 83.75 | ||||||||||
-| SICK-R | 82.38 | 89.62 | 78.96 | 78.63 | 75.83 | 78.95 | ||||||||||
-| STS12 | 76.72 | 82.43 | 73.41 | 74.57 | 73.60 | 73.52 | ||||||||||
-| STS13 | 79.90 | 89.49 | 80.25 | 82.35 | 76.50 | 83.24 | ||||||||||
-| STS14 | 77.49 | 84.95 | 75.13 | 76.89 | 76.35 | 78.68 | ||||||||||
-| STS15 | 86.42 | 89.31 | 84.98 | 85.62 | 80.90 | 87.46 | ||||||||||
-| STS16 | 85.18 | 86.85 | 83.70 | 82.66 | 74.73 | 83.77 | ||||||||||
-| STS17 | 87.87 | 89.99 | 86.13 | 86.95 | 83.84 | 89.77 | ||||||||||
-| STS22 | 67.07 | 67.28 | 66.31 | 67.58 | 62.52 | 65.15 | ||||||||||
-| STSBenchmark | 86.87 | 89.44 | 84.27 | 84.40 | 79.64 | 84.93 | ||||||||||
-| Task |
-| Retrieval - nDCG@5 | |||||||||||||
-| Average | 52.65 | 37.37 | 47.76 | 17.72 | 46.55 | ||||||||
-| ArxivQ | 64.92 | 58.78 | 66.14 | 25.40 | 30.44 | ||||||||
-| DocQ | 24.64 | 21.16 | 24.81 | 11.90 | 23.82 | ||||||||
-| InfoQ | 57.90 | 56.62 | 56.63 | 35.50 | 59.87 | ||||||||
-| TabF | 45.91 | 37.28 | 43.18 | 20.20 | 70.68 | ||||||||
-| TATQ | 30.25 | 12.09 | 25.79 | 3.30 | 20.00 | ||||||||
-| Shift | 34.07 | 28.41 | 31.54 | 3.80 | 30.79 | ||||||||
-| AI | 68.07 | 32.95 | 55.64 | 15.20 | 47.90 | ||||||||
-| Energy | 62.15 | 49.15 | 60.97 | 19.70 | 64.94 | ||||||||
-| Gov | 68.97 | 37.81 | 55.56 | 21.40 | 58.62 | ||||||||
-| Health | 69.05 | 39.50 | 57.37 | 20.80 | 58.43 | ||||||||
+| Zero-shot Image Retrieval - Recall@5 [%] | ||||||||||||||||
+| average | 84.87 | 80.66 | 86.11 | 31.01 | 87.60 | 86.23 | ||||||||||
+| de | 85.70 | 80.40 | 86.80 | 48.80 | 88.30 | 87.00 | ||||||||||
+| en | 89.40 | 84.00 | 89.40 | 89.00 | 89.40 | 88.80 | ||||||||||
+| es | 85.90 | 82.70 | 87.40 | 56.80 | 88.20 | 86.70 | ||||||||||
+| fr | 85.10 | 80.90 | 87.30 | 66.80 | 87.70 | 86.90 | ||||||||||
+| it | 85.80 | 83.20 | 87.10 | 45.00 | 89.30 | 87.80 | ||||||||||
+| ko | 82.10 | 78.30 | 83.00 | 1.30 | 85.20 | 83.60 | ||||||||||
+| pl | 86.50 | 81.10 | 88.00 | 11.00 | 89.40 | 87.60 | ||||||||||
+| ru | 81.10 | 79.30 | 82.10 | 3.20 | 83.40 | 82.40 | ||||||||||
+| tr | 83.70 | 81.30 | 86.30 | 7.00 | 88.30 | 86.80 | ||||||||||
+| zh | 83.40 | 75.40 | 83.70 | 4.50 | 86.80 | 84.70 | ||||||||||
+| Zero-shot Text Retrieval - Recall@5 [%] | ||||||||||||||||
+| average | 86.03 | 83.02 | 86.45 | 36.89 | 85.37 | 84.56 | ||||||||||
+| de | 86.20 | 82.90 | 86.90 | 50.10 | 84.40 | 84.30 | ||||||||||
+| en | 90.50 | 87.80 | 90.60 | 89.40 | 88.30 | 87.50 | ||||||||||
+| es | 87.00 | 84.90 | 87.50 | 55.60 | 85.80 | 85.20 | ||||||||||
+| fr | 85.20 | 82.20 | 85.00 | 67.30 | 86.30 | 84.40 | ||||||||||
+| it | 88.00 | 83.10 | 87.70 | 51.10 | 86.70 | 85.60 | ||||||||||
+| ko | 82.10 | 79.30 | 84.00 | 1.60 | 83.20 | 82.60 | ||||||||||
+| pl | 88.50 | 84.50 | 88.90 | 15.50 | 86.80 | 86.60 | ||||||||||
+| ru | 83.10 | 79.60 | 81.80 | 3.50 | 80.90 | 81.20 | ||||||||||
+| tr | 85.60 | 83.20 | 85.90 | 9.50 | 87.10 | 85.40 | ||||||||||
+| zh | 84.10 | 82.70 | 86.20 | 7.40 | 84.20 | 82.80 | ||||||||||
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| AILACasedocs | test | 32.79 | 6.62 | 8.63 | 34.73 | 29.21 | 31.88 | 23.10 | ||
+| AILAStatutes | test | 14.45 | 15.04 | 13.58 | 33.00 | 22.58 | 26.45 | 8.86 | ||
+| ARCChallenge | test | 10.52 | 5.85 | 6.78 | 10.12 | 11.18 | 10.92 | 10.96 | ||
+| AlphaNLI | test | 31.45 | 16.02 | 19.25 | 30.16 | 28.39 | 27.87 | 29.12 | ||
+| ArguAna | test | 49.41 | 25.01 | 33.81 | 43.28 | 42.85 | 47.99 | 43.65 | ||
+| BelebeleRetrieval | test | 91.83 | 72.88 | 77.43 | 93.25 | 91.84 | 91.86 | 92.97 | ||
+test | 52.14 | 27.46 | 31.30 | 37.01 | 52.14 | 52.58 | 50.21 | ||
+test | 45.74 | 19.98 | 23.40 | 44.50 | 48.67 | 48.19 | 46.36 | ||
+test | 59.61 | 34.23 | 37.64 | 47.53 | 58.56 | 58.39 | 54.40 | ||
+test | 39.30 | 17.70 | 20.62 | 35.66 | 38.72 | 39.48 | 40.61 | ||
+test | 28.23 | 13.08 | 15.66 | 30.23 | 30.69 | 31.37 | 30.70 | ||
+test | 46.31 | 25.37 | 28.74 | 42.18 | 45.62 | 46.61 | 45.84 | ||
+test | 41.40 | 21.06 | 25.66 | 37.33 | 41.32 | 42.61 | 41.30 | ||
+| CQADupstackRetrieval | test | 40.93 | 20.74 | 23.54 | 36.02 | 41.48 | 42.17 | 41.19 | ||
+test | 33.97 | 19.10 | 20.63 | 31.20 | 35.06 | 35.43 | 35.41 | ||
+test | 29.43 | 12.42 | 13.46 | 26.97 | 29.71 | 30.41 | 30.92 | ||
+test | 41.13 | 19.57 | 20.99 | 37.22 | 42.60 | 43.01 | 42.48 | ||
+test | 41.43 | 23.31 | 26.58 | 33.16 | 41.81 | 43.94 | 42.01 | ||
+test | 32.43 | 15.65 | 17.87 | 29.20 | 32.92 | 34.05 | 34.05 | ||
+| CUREv1 | dentistry_and_oral_health | 45.53 | 20.08 | 26.11 | 49.19 | 41.41 | 40.88 | 50.26 | ||
+| CUREv1 | dermatology | 51.32 | 17.25 | 22.31 | 53.54 | 41.41 | 42.34 | 59.59 | ||
+| CUREv1 | gastroenterology | 45.20 | 19.99 | 23.25 | 49.36 | 40.56 | 41.59 | 51.96 | ||
+| CUREv1 | genetics | 48.26 | 18.91 | 27.19 | 56.64 | 41.62 | 43.95 | 54.23 | ||
+| CUREv1 | neuroscience_and_neurology | 38.67 | 15.97 | 21.31 | 44.91 | 35.35 | 35.65 | 42.92 | ||
+| CUREv1 | orthopedic_surgery | 45.13 | 12.79 | 19.48 | 44.11 | 40.72 | 38.91 | 46.29 | ||
+| CUREv1 | otorhinolaryngology | 39.65 | 12.28 | 18.19 | 41.91 | 34.29 | 32.90 | 46.59 | ||
+| CUREv1 | plastic_surgery | 46.10 | 15.43 | 23.43 | 46.86 | 40.10 | 40.62 | 48.34 | ||
+| CUREv1 | psychiatry_and_psychology | 46.87 | 21.02 | 27.88 | 52.34 | 41.78 | 42.10 | 50.63 | ||
+| CUREv1 | pulmonology | 45.08 | 16.75 | 26.42 | 53.72 | 41.09 | 40.11 | 51.80 | ||
+| CUREv1 | avg | 45.18 | 17.05 | 23.56 | 49.26 | 39.83 | 39.91 | 50.26 | ||
+test | 78.05 | 41.81 | 44.20 | 68.65 | 63.21 | 61.77 | 58.96 | ||
+| ChemNQRetrieval | test | 57.11 | 39.22 | 39.19 | 60.80 | 48.21 | 47.78 | 54.58 | ||
+| ClimateFEVER | test | 24.82 | 13.06 | 16.44 | 42.26 | 19.31 | 23.70 | 30.14 | ||
+test | 25.32 | 14.03 | 17.11 | 43.02 | 20.16 | 24.22 | 30.47 | ||
+| CodeFeedbackMT | test | 39.15 | 4.86 | 6.66 | 59.84 | 50.42 | 52.46 | 47.11 | ||
+| CodeFeedbackST | test | 66.06 | 17.09 | 20.64 | 78.09 | 70.81 | 71.05 | 68.62 | ||
+| DBPedia | test | 36.66 | 17.76 | 21.37 | 41.00 | 22.03 | 24.55 | 37.42 | ||
+| DBPediaHardNegatives | test | 38.84 | 23.32 | 26.66 | 43.64 | 26.47 | 28.54 | 40.76 | ||
+| FEVER | test | 76.30 | 29.84 | 33.91 | 89.07 | 60.54 | 67.09 | 84.98 | ||
+| FEVERHardNegatives | test | 77.02 | 38.27 | 40.69 | 89.86 | 61.52 | 68.72 | 86.56 | ||
+| FaithDial | test | 23.18 | 16.83 | 18.69 | 26.35 | 24.03 | 23.93 | 22.84 | ||
+| FeedbackQARetrieval | test | 49.15 | 22.69 | 33.43 | 52.06 | 45.38 | 49.90 | 51.00 | ||
+| FiQA2018 | test | 38.27 | 8.74 | 12.79 | 47.46 | 41.65 | 40.96 | 42.00 | ||
+| HagridRetrieval | dev | 98.69 | 96.85 | 97.16 | 98.69 | 98.65 | 98.42 | 98.69 | ||
+| HellaSwag | test | 27.30 | 14.33 | 18.39 | 29.26 | 29.32 | 29.02 | 27.21 | ||
+| HotpotQA | test | 61.89 | 21.55 | 23.09 | 64.65 | 49.63 | 46.96 | 60.15 | ||
+test | 62.58 | 27.56 | 28.59 | 64.74 | 52.00 | 49.67 | 60.01 | ||
+test | 33.55 | 10.82 | 12.62 | 34.22 | 19.86 | 19.68 | 23.57 | ||
+| LEMBNeedleRetrieval | test_256 | 72.00 | 36.00 | 40.00 | 64.00 | 78.00 | 72.00 | 86.00 | ||
+| LEMBNeedleRetrieval | test_512 | 58.00 | 24.00 | 18.00 | 32.00 | 64.00 | 54.00 | 70.00 | ||
+| LEMBNeedleRetrieval | test_1024 | 66.00 | 6.00 | 6.00 | 14.00 | 64.00 | 50.00 | 46.00 | ||
+| LEMBNeedleRetrieval | test_2048 | 76.00 | 8.00 | 8.00 | 10.00 | 60.00 | 56.00 | 58.00 | ||
+| LEMBNeedleRetrieval | test_4096 | 66.00 | 2.00 | 2.00 | 8.00 | 48.00 | 48.00 | 54.00 | ||
+| LEMBNeedleRetrieval | test_8192 | 72.00 | 0.00 | 2.00 | 12.00 | 46.00 | 46.00 | 34.00 | ||
+| LEMBNeedleRetrieval | test_16384 | 42.00 | 0.00 | 0.00 | 8.00 | 28.00 | 26.00 | 28.00 | ||
+| LEMBNeedleRetrieval | test_32768 | 16.00 | 2.00 | 4.00 | 2.00 | 6.00 | 8.00 | 8.00 | ||
+| LEMBNeedleRetrieval | avg | 58.50 | 9.75 | 10.00 | 18.75 | 49.25 | 45.00 | 48.00 | ||
+| LEMBPasskeyRetrieval | test_256 | 90.00 | 22.00 | 24.00 | 100.00 | 98.00 | 76.00 | 100.00 | ||
+| LEMBPasskeyRetrieval | test_512 | 52.00 | 20.00 | 16.00 | 100.00 | 58.00 | 72.00 | 98.00 | ||
+| LEMBPasskeyRetrieval | test_1024 | 34.00 | 14.00 | 14.00 | 94.00 | 16.00 | 42.00 | 56.00 | ||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| LEMBPasskeyRetrieval | test_2048 | 34.00 | 4.00 | 4.00 | 72.00 | 12.00 | 14.00 | 12.00 | ||
+| LEMBPasskeyRetrieval | test_4096 | 72.00 | 0.00 | 0.00 | 42.00 | 2.00 | 20.00 | 0.00 | ||
+| LEMBPasskeyRetrieval | test_8192 | 82.00 | 0.00 | 2.00 | 26.00 | 10.00 | 4.00 | 8.00 | ||
+| LEMBPasskeyRetrieval | test_16384 | 44.00 | 2.00 | 2.00 | 26.00 | 8.00 | 4.00 | 6.00 | ||
+| LEMBPasskeyRetrieval | test_32768 | 26.00 | 2.00 | 0.00 | 10.00 | 6.00 | 8.00 | 4.00 | ||
+| LEMBPasskeyRetrieval | avg | 54.25 | 8.00 | 7.75 | 58.75 | 26.25 | 30.00 | 35.50 | ||
+| LEMBQMSumRetrieval | test | 37.38 | 8.05 | 9.57 | 39.35 | 29.37 | 29.93 | 31.07 | ||
+validation | 93.37 | 23.44 | 23.31 | 92.22 | 55.86 | 56.84 | 75.02 | ||
+| LEMBWikimQARetrieval | test | 75.19 | 33.00 | 35.49 | 65.95 | 57.39 | 55.88 | 64.85 | ||
+test | 69.55 | 39.91 | 44.37 | 78.14 | 61.34 | 67.18 | 64.18 | ||
+test | 89.09 | 81.55 | 86.17 | 93.65 | 88.40 | 90.37 | 91.57 | ||
+| LegalSummarization | test | 63.23 | 45.25 | 50.57 | 59.25 | 53.08 | 56.06 | 62.70 | ||
+| LitSearchRetrieval | test | 41.89 | 11.96 | 12.94 | 48.18 | 39.51 | 39.15 | 48.37 | ||
+dev | 43.34 | 19.57 | 20.44 | 51.97 | 31.97 | 39.01 | 50.77 | ||
+| MLQARetrieval | test | 63.42 | 35.82 | 40.61 | 64.72 | 59.87 | 60.03 | 64.08 | ||
+| MLQuestions | test | 59.43 | 27.79 | 30.23 | 63.10 | 54.91 | 54.74 | 56.11 | ||
+| MSMARCO | dev | 36.92 | 10.63 | 13.80 | 40.84 | 22.94 | 27.12 | 37.40 | ||
+| MSMARCOHardNegatives | test | 65.43 | 41.81 | 44.71 | 71.74 | 55.21 | 58.27 | 69.89 | ||
+| MedicalQARetrieval | test | 67.24 | 34.04 | 39.22 | 70.23 | 66.01 | 68.24 | 71.34 | ||
+test | 34.78 | 8.38 | 10.80 | 28.87 | 19.50 | 18.97 | 20.16 | ||
+| NFCorpus | test | 33.52 | 16.77 | 22.13 | 36.61 | 31.00 | 32.16 | 32.88 | ||
+| NQ | test | 58.11 | 16.28 | 19.94 | 64.33 | 40.00 | 43.03 | 57.16 | ||
+| NQHardNegatives | test | 59.43 | 21.05 | 23.33 | 65.06 | 41.42 | 44.50 | 58.50 | ||
+| NanoArguAnaRetrieval | train | 55.13 | 30.90 | 37.50 | 49.81 | 49.93 | 59.43 | 51.27 | ||
+train | 29.09 | 21.87 | 27.72 | 44.57 | 28.17 | 29.40 | 34.37 | ||
+| NanoDBPediaRetrieval | train | 57.42 | 40.98 | 41.53 | 61.81 | 49.16 | 53.60 | 59.45 | ||
+| NanoFEVERRetrieval | train | 88.59 | 56.74 | 57.84 | 91.93 | 80.95 | 82.60 | 90.54 | ||
+train | 46.25 | 17.92 | 26.05 | 53.61 | 53.80 | 53.22 | 53.41 | ||
+train | 75.15 | 42.22 | 51.20 | 75.65 | 67.68 | 65.87 | 72.86 | ||
+| NanoMSMARCORetrieval | train | 61.17 | 30.93 | 41.23 | 66.53 | 48.93 | 53.50 | 61.97 | ||
+train | 34.05 | 16.33 | 20.63 | 36.85 | 29.81 | 31.01 | 32.01 | ||
+| NanoNQRetrieval | train | 69.86 | 28.61 | 36.17 | 71.94 | 61.82 | 57.89 | 66.80 | ||
+| NanoQuoraRetrieval | train | 94.66 | 92.18 | 89.85 | 76.55 | 94.22 | 93.80 | 95.87 | ||
+| NanoSCIDOCSRetrieval | train | 38.47 | 21.36 | 28.36 | 40.42 | 41.56 | 41.27 | 39.07 | ||
+| NanoSciFactRetrieval | train | 71.75 | 24.97 | 33.39 | 79.08 | 71.41 | 72.98 | 70.20 | ||
+train | 50.02 | 30.00 | 36.17 | 53.24 | 45.07 | 48.62 | 52.28 | ||
+| NarrativeQARetrieval | test | 33.45 | 11.31 | 12.58 | 34.13 | 19.83 | 19.63 | 23.51 | ||
+| PIQA | test | 30.40 | 12.58 | 15.64 | 32.02 | 28.67 | 30.35 | 30.65 | ||
+| PublicHealthQA | test | 79.51 | 56.80 | 64.25 | 83.56 | 81.30 | 84.56 | 84.12 | ||
+| Quail | test | 4.38 | 1.51 | 1.85 | 4.20 | 4.11 | 4.01 | 3.33 | ||
+| QuoraRetrieval | test | 87.88 | 77.12 | 80.47 | 61.68 | 87.14 | 86.98 | 88.14 | ||
+test | 87.88 | 77.08 | 80.41 | 63.39 | 87.39 | 87.24 | 87.98 | ||
+| RARbCode | test | 38.02 | 0.21 | 1.77 | 54.61 | 42.50 | 41.59 | 35.27 | ||
+| RARbMath | test | 53.29 | 14.87 | 18.83 | 74.57 | 61.83 | 62.81 | 59.99 | ||
+| SCIDOCS | test | 20.23 | 9.87 | 10.87 | 19.91 | 20.51 | 20.85 | 18.93 | ||
+| SCIDOCS-NL | test | 8.19 | 5.30 | 7.94 | 17.16 | 14.43 | 16.71 | 15.50 | ||
+| SIQA | test | 1.87 | 0.81 | 0.76 | 0.79 | 1.99 | 2.27 | 2.31 | ||
+| SciFact | test | 67.32 | 28.65 | 33.64 | 72.56 | 68.18 | 66.95 | 65.21 | ||
+| SpartQA | test | 7.67 | 0.67 | 3.22 | 0.73 | 0.10 | 1.82 | 8.51 | ||
+| StackOverflowQA | test | 82.28 | 14.30 | 19.87 | 90.79 | 81.33 | 86.46 | 84.43 | ||
+test | 26.33 | 14.73 | 15.49 | 33.44 | 4.26 | 11.41 | 26.09 | ||
+| TRECCOVID | test | 71.58 | 37.48 | 45.05 | 77.34 | 51.51 | 58.65 | 76.69 | ||
+| TempReasonL1 | test | 1.49 | 0.27 | 0.44 | 0.60 | 1.12 | 1.19 | 1.19 | ||
+| TempReasonL2Context | test | 9.15 | 5.95 | 7.88 | 11.73 | 6.85 | 6.78 | 7.82 | ||
+| TempReasonL2Fact | test | 15.07 | 6.26 | 7.92 | 19.70 | 11.98 | 11.72 | 16.03 | ||
+| TempReasonL2Pure | test | 0.95 | 0.12 | 0.23 | 0.50 | 0.96 | 1.14 | 1.03 | ||
+| TempReasonL3Context | test | 8.73 | 6.78 | 7.79 | 11.62 | 5.36 | 5.70 | 6.81 | ||
+| TempReasonL3Fact | test | 13.43 | 6.94 | 7.29 | 18.35 | 9.69 | 10.19 | 12.33 | ||
+| TempReasonL3Pure | test | 6.30 | 3.62 | 3.86 | 5.45 | 3.65 | 4.70 | 4.50 | ||
+| TopiOCQA | validation | 14.13 | 5.59 | 6.04 | 19.18 | 12.52 | 12.04 | 15.16 | ||
+validation | 14.02 | 6.17 | 6.98 | 19.19 | 13.23 | 12.58 | 15.03 |
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+test | 57.00 | 27.64 | 33.36 | 55.40 | 47.28 | 52.60 | 56.19 | ||
+test | 90.91 | 73.24 | 76.80 | 91.78 | 89.89 | 90.40 | 91.85 | ||
+| WinoGrande | test | 49.60 | 45.80 | 44.84 | 19.58 | 22.21 | 28.12 | 43.78 | ||
+| XMarket | test | 30.73 | 7.26 | 13.61 | 34.79 | 14.79 | 21.83 | 29.33 | ||
+| XQuADRetrieval | validation | 94.73 | 75.69 | 79.50 | 96.11 | 92.72 | 92.97 | 94.79 | ||
+| Average | 47.04 | 24.69 | 27.83 | 48.34 | 42.02 | 43.47 | 46.46 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| BIOSSES | test | 83.75 | 65.77 | 60.63 | 84.52 | 83.53 | 83.14 | 82.90 | ||
+| SICK-R | test | 78.95 | 73.73 | 75.82 | 77.25 | 78.97 | 78.72 | 82.40 | ||
+| STS12 | test | 73.52 | 70.67 | 73.33 | 78.34 | 73.43 | 74.52 | 76.71 | ||
+| STS13 | test | 83.24 | 73.13 | 76.43 | 87.10 | 80.22 | 82.29 | 79.92 | ||
+| STS14 | test | 78.67 | 69.76 | 72.23 | 80.15 | 75.11 | 76.88 | 77.50 | ||
+| STS15 | test | 87.46 | 80.87 | 80.90 | 87.24 | 84.99 | 85.57 | 86.43 | ||
+| STS16 | test | 83.77 | 73.88 | 77.03 | 83.83 | 83.71 | 82.71 | 85.19 | ||
+| STS17 | test | 89.78 | 82.77 | 83.66 | 88.33 | 87.89 | 86.93 | 87.92 | ||
+| STS22.v2 | test | 65.84 | 55.21 | 58.06 | 65.50 | 63.46 | 68.19 | 67.00 | ||
+| STSBenchmark | test | 84.93 | 74.51 | 79.45 | 85.09 | 84.30 | 84.46 | 86.86 | ||
+test | 84.93 | 74.51 | 79.45 | 85.09 | 84.30 | 84.45 | 86.86 | ||
+| SemRel24STS | test | 81.36 | 76.07 | 79.75 | 77.70 | 80.19 | 81.29 | 79.31 | ||
+| Average | 81.35 | 72.57 | 74.73 | 81.68 | 80.01 | 80.76 | 81.58 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 15.96 | 63.34 | 70.18 | 91.32 | 92.23 | 92.47 | 91.57 | ||
+| BelebeleRetrieval | test | 17.04 | 62.98 | 68.77 | 92.24 | 91.50 | 91.16 | 90.81 | ||
+| CmedqaRetrieval | dev | 1.81 | 5.09 | 7.42 | 35.89 | 32.48 | 31.98 | 31.73 | ||
+| CovidRetrieval | dev | 1.37 | 21.33 | 23.15 | 78.91 | 66.57 | 68.60 | 72.18 | ||
+| DuRetrieval | dev | 3.92 | 19.05 | 27.65 | 83.11 | 76.88 | 77.85 | 78.50 | ||
+| EcomRetrieval | dev | 4.33 | 13.65 | 22.54 | 60.68 | 29.77 | 47.77 | 55.86 | ||
+| LeCaRDv2 | test | 23.50 | 21.92 | 30.11 | 58.40 | 55.44 | 54.00 | 50.64 | ||
+dev | 0.34 | 16.15 | 18.26 | 57.66 | 30.84 | 40.22 | 55.62 | ||
+| MLQARetrieval | test | 6.29 | 31.77 | 36.78 | 64.72 | 58.95 | 59.40 | 64.69 | ||
+| MMarcoRetrieval | dev | 7.06 | 28.30 | 37.94 | 79.66 | 70.05 | 69.96 | 76.28 | ||
+| MedicalRetrieval | dev | 1.32 | 7.73 | 11.96 | 56.63 | 52.92 | 51.84 | 51.15 | ||
+test | 0.54 | 3.32 | 3.51 | 17.17 | 10.79 | 11.37 | 7.78 | ||
+test | 1.99 | 24.86 | 29.41 | 54.55 | 39.14 | 44.28 | 49.03 | ||
+test | 2.50 | 24.44 | 30.14 | 50.10 | 42.90 | 43.89 | 49.84 | ||
+| PublicHealthQA | test | 17.66 | 59.55 | 65.23 | 84.56 | 85.38 | 86.30 | 84.94 | ||
+| T2Retrieval | dev | 2.77 | 21.45 | 28.94 | 83.16 | 72.88 | 76.19 | 77.76 | ||
+| VideoRetrieval | dev | 6.14 | 17.77 | 24.01 | 70.45 | 30.10 | 42.45 | 59.71 | ||
+| XPQARetrieval | test | 18.03 | 30.56 | 42.41 | 69.54 | 66.79 | 66.23 | 65.72 | ||
+| XQuADRetrieval | validation | 28.53 | 69.93 | 74.71 | 93.83 | 92.34 | 92.77 | 94.08 | ||
+test | 0.61 | -2.44 | 6.03 | 2.80 | 4.33 | 4.39 | 1.57 | ||
+| Average | 8.09 | 27.04 | 32.96 | 64.27 | 55.11 | 57.66 | 60.47 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| AFQMC | validation | 7.65 | 10.97 | 14.60 | 38.85 | 35.27 | 36.19 | 36.56 | ||
+| ATEC | test | 14.16 | 13.99 | 21.44 | 44.80 | 42.25 | 42.87 | 43.59 | ||
+| BQ | test | 22.52 | 28.15 | 33.65 | 47.27 | 46.78 | 45.86 | 55.03 | ||
+| LCQMC | test | 20.96 | 40.26 | 52.80 | 74.47 | 73.56 | 74.16 | 75.42 | ||
+| PAWSX | test | 8.29 | 10.63 | 13.00 | 15.26 | 14.20 | 15.96 | 15.51 | ||
+| QBQTC | test | 17.95 | 22.54 | 22.54 | 34.35 | 30.09 | 32.60 | 32.74 | ||
+| STS22.v2 | test | 44.52 | 43.00 | 50.37 | 72.65 | 71.58 | 71.00 | 71.38 | ||
+| STSB | test | 34.50 | 65.25 | 72.05 | 81.35 | 80.18 | 80.52 | 82.70 | ||
+test | 36.19 | 65.46 | 71.87 | 82.42 | 80.78 | 81.15 | 82.75 | ||
+| Average | 22.97 | 33.36 | 39.15 | 54.60 | 52.74 | 53.37 | 55.08 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 3.61 | 63.82 | 69.57 | 89.21 | 86.52 | 87.21 | 86.21 | ||
+| BelebeleRetrieval | test | 43.80 | 28.07 | 31.67 | 66.80 | 66.14 | 64.76 | 62.00 | ||
+| IndicQARetrieval | test | 2.11 | 32.46 | 35.63 | 67.26 | 63.22 | 62.60 | 63.87 | ||
+dev | 0.14 | 21.87 | 23.79 | 56.47 | 43.20 | 43.77 | 54.80 | ||
+| MLQARetrieval | test | 1.82 | 34.81 | 40.69 | 63.77 | 59.08 | 58.41 | 62.40 | ||
+| MintakaRetrieval | test | 1.89 | 19.62 | 23.41 | 24.42 | 23.87 | 24.72 | 26.48 | ||
+test | 1.49 | 8.68 | 10.52 | 25.40 | 26.93 | 30.47 | 23.68 | ||
+test | 3.82 | 54.61 | 57.96 | 86.57 | 83.89 | 82.75 | 85.09 | ||
+| XPQARetrieval | test | 20.45 | 57.18 | 64.98 | 78.26 | 77.26 | 77.50 | 74.36 | ||
+| XQuADRetrieval | validation | 8.21 | 70.70 | 74.00 | 93.66 | 90.16 | 89.68 | 91.90 | ||
+| Average | 8.73 | 39.18 | 43.22 | 65.18 | 62.03 | 62.19 | 63.08 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| SemRel24STS | test | 36.09 | 68.11 | 73.78 | 83.62 | 83.35 | 84.30 | 79.55 | ||
+| Average | 36.09 | 68.11 | 73.78 | 83.62 | 83.35 | 84.30 | 79.55 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 54.13 | 39.84 | 59.47 | 92.99 | 93.20 | 93.21 | 92.08 | ||
+| GerDaLIR | test | 2.14 | 0.14 | 0.41 | 16.18 | 9.06 | 10.57 | 2.22 | ||
+| GerDaLIRSmall | test | 6.18 | 0.47 | 1.34 | 36.66 | 21.23 | 25.33 | 6.76 | ||
+| GermanDPR | test | 49.26 | 35.74 | 45.63 | 82.47 | 80.26 | 81.00 | 80.52 | ||
+test | 45.22 | 37.40 | 47.70 | 88.71 | 83.07 | 82.75 | 87.18 | ||
+| GermanQuAD-Retrieval | test | 62.67 | 47.35 | 61.08 | 94.15 | 90.77 | 91.66 | 92.80 | ||
+| LegalQuAD | test | 14.97 | 4.06 | 11.18 | 58.83 | 41.33 | 46.25 | 32.87 | ||
+dev | 19.37 | 13.07 | 20.59 | 53.37 | 39.41 | 42.67 | 51.61 | ||
+| MLQARetrieval | test | 33.09 | 17.87 | 28.39 | 67.50 | 63.06 | 61.94 | 65.77 | ||
+| MintakaRetrieval | test | 20.37 | 9.66 | 17.62 | 27.05 | 29.08 | 30.10 | 31.55 | ||
+test | 13.25 | 3.15 | 9.74 | 37.81 | 34.15 | 38.18 | 24.10 | ||
+test | 62.29 | 44.00 | 58.90 | 89.34 | 89.01 | 88.04 | 88.75 | ||
+| XMarket | test | 8.78 | 4.86 | 7.76 | 28.28 | 12.12 | 18.33 | 16.54 | ||
+| XPQARetrieval | test | 52.72 | 32.65 | 48.28 | 84.68 | 83.44 | 83.50 | 82.31 | ||
+| XQuADRetrieval | validation | 62.08 | 49.82 | 66.27 | 95.79 | 94.18 | 93.79 | 94.90 | ||
+| Average | 33.77 | 22.67 | 32.29 | 63.59 | 57.56 | 59.16 | 56.66 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| GermanSTSBenchmark | test | 63.94 | 54.53 | 66.52 | 82.98 | 82.17 | 82.25 | 84.82 | ||
+| STS22.v2 | test | 22.42 | 10.39 | 27.08 | 59.16 | 59.22 | 61.35 | 58.30 | ||
+test | 63.78 | 53.82 | 66.77 | 83.83 | 83.14 | 83.33 | 85.68 | ||
+| Average | 50.04 | 39.58 | 53.45 | 75.32 | 74.84 | 75.64 | 76.27 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| AlloprofRetrieval | test | 22.90 | 12.36 | 18.70 | 54.39 | 47.52 | 51.75 | 51.24 | ||
+| BSARDRetrieval | test | 25.68 | 19.37 | 36.49 | 63.96 | 64.86 | 65.77 | 58.11 | ||
+| BelebeleRetrieval | test | 66.86 | 53.05 | 71.63 | 93.86 | 90.26 | 90.65 | 92.38 | ||
+| FQuADRetrieval | test | 51.07 | 36.70 | 43.68 | 74.33 | 68.04 | 67.03 | 70.59 | ||
+dev | 26.57 | 17.58 | 19.93 | 55.22 | 38.43 | 42.81 | 52.72 | ||
+| MintakaRetrieval | test | 22.91 | 15.57 | 20.15 | 26.94 | 28.39 | 29.17 | 30.79 | ||
+test | 41.80 | 20.64 | 26.17 | 59.84 | 59.76 | 66.51 | 54.17 | ||
+| PublicHealthQA | test | 64.54 | 51.62 | 64.67 | 92.05 | 88.13 | 90.71 | 91.75 | ||
+test | 2.31 | 5.09 | 11.42 | 22.68 | 4.63 | 10.65 | 17.13 | ||
+| SyntecRetrieval | test | 64.17 | 44.47 | 63.52 | 84.03 | 77.33 | 78.95 | 79.01 | ||
+| XPQARetrieval | test | 51.10 | 34.80 | 49.81 | 77.68 | 76.80 | 77.56 | 74.84 | ||
+| Average | 39.99 | 28.30 | 38.74 | 64.09 | 58.56 | 61.05 | 61.16 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| SICKFr | test | 67.42 | 63.22 | 68.86 | 76.51 | 77.38 | 77.02 | 80.55 | ||
+| STS22.v2 | test | 66.65 | 61.01 | 73.31 | 83.48 | 77.99 | 81.16 | 79.90 | ||
+test | 69.82 | 67.60 | 74.97 | 83.33 | 81.79 | 82.40 | 84.65 | ||
+| Average | 67.97 | 63.94 | 72.38 | 81.11 | 79.06 | 80.19 | 81.70 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 62.63 | 52.86 | 69.01 | 93.52 | 91.81 | 91.42 | 92.36 | ||
+dev | 27.05 | 16.71 | 22.04 | 51.08 | 43.17 | 45.33 | 50.94 | ||
+| MLQARetrieval | test | 42.28 | 29.25 | 40.56 | 68.62 | 62.89 | 61.50 | 66.69 | ||
+| MintakaRetrieval | test | 21.47 | 16.29 | 20.21 | 26.93 | 29.00 | 29.76 | 31.28 | ||
+test | 41.13 | 14.04 | 20.55 | 62.07 | 63.86 | 67.68 | 55.56 | ||
+| PublicHealthQA | test | 53.89 | 46.09 | 61.15 | 83.09 | 83.61 | 86.62 | 83.75 | ||
+test | 20.02 | 18.03 | 28.32 | 43.11 | 35.95 | 38.94 | 41.72 | ||
+test | 38.77 | 37.06 | 48.75 | 69.71 | 70.12 | 70.79 | 72.98 | ||
+| XMarket | test | 12.02 | 6.38 | 10.22 | 26.69 | 14.20 | 20.96 | 17.51 | ||
+| XPQARetrieval | test | 44.80 | 29.19 | 42.81 | 72.14 | 71.04 | 71.36 | 69.10 | ||
+| XQuADRetrieval | validation | 73.39 | 64.76 | 75.47 | 95.67 | 92.84 | 92.20 | 94.40 | ||
+| Average | 39.77 | 30.06 | 39.92 | 62.97 | 59.86 | 61.51 | 61.48 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| STS17 | test | 77.09 | 77.23 | 83.26 | 86.93 | 87.31 | 86.99 | 87.29 | ||
+| STS22.v2 | test | 57.19 | 46.44 | 53.42 | 75.58 | 75.16 | 76.48 | 76.01 | ||
+test | 71.90 | 65.44 | 75.15 | 84.72 | 82.76 | 83.35 | 85.90 | ||
+| STSES | test | 74.00 | 72.02 | 66.90 | 75.89 | 76.47 | 75.22 | 78.34 | ||
+| Average | 70.05 | 65.28 | 69.68 | 80.78 | 80.42 | 80.51 | 81.89 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 17.69 | 64.01 | 69.81 | 91.73 | 91.12 | 90.84 | 90.23 | ||
+| JaGovFaqsRetrieval | test | 8.85 | 34.49 | 40.39 | 71.92 | 72.73 | 72.10 | 70.31 | ||
+| JaQuADRetrieval | validation | 8.62 | 29.17 | 32.74 | 54.78 | 51.98 | 53.23 | 52.43 | ||
+| JaqketRetrieval | test | 0.09 | 19.51 | 26.41 | 46.91 | 25.71 | 26.54 | 32.70 | ||
+dev | 0.82 | 17.52 | 18.26 | 66.26 | 47.37 | 50.65 | 57.92 | ||
+| MintakaRetrieval | test | 5.38 | 15.54 | 17.89 | 20.68 | 24.43 | 24.77 | 25.57 | ||
+test | 0.49 | 8.79 | 10.37 | 38.40 | 28.15 | 38.88 | 33.15 | ||
+test | 22.23 | 36.99 | 37.98 | 98.75 | 96.25 | 97.28 | 97.78 | ||
+test | 17.13 | 49.47 | 56.92 | 94.73 | 95.53 | 95.25 | 95.77 | ||
+test | 4.95 | 19.20 | 23.94 | 93.40 | 87.58 | 90.32 | 91.21 | ||
+| XPQARetrieval | test | 32.25 | 43.38 | 56.12 | 75.55 | 75.89 | 76.56 | 73.22 | ||
+| Average | 10.77 | 30.73 | 35.53 | 68.47 | 63.34 | 65.13 | 65.48 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| JSICK | test | 60.57 | 77.93 | 78.38 | 81.38 | 80.56 | 80.28 | 80.45 | ||
+| JSTS | validation | 51.09 | 76.85 | 80.11 | 80.69 | 82.39 | 82.64 | 83.57 | ||
+| Average | 55.83 | 77.39 | 79.25 | 81.03 | 81.48 | 81.46 | 82.01 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| BelebeleRetrieval | test | 9.30 | 63.01 | 71.61 | 93.11 | 90.74 | 90.96 | 92.41 | ||
+dev | 1.15 | 18.44 | 21.74 | 65.29 | 42.54 | 49.71 | 59.74 | ||
+test | 2.85 | 15.13 | 20.25 | 49.54 | 54.01 | 58.75 | 46.55 | ||
+test | 0.97 | 29.00 | 36.59 | 58.08 | 47.32 | 47.36 | 52.48 | ||
+test | 4.26 | 28.95 | 34.23 | 52.91 | 47.55 | 49.54 | 51.30 | ||
+| PublicHealthQA | test | 16.70 | 55.78 | 60.23 | 84.49 | 85.86 | 87.28 | 84.79 | ||
+| RiaNewsRetrieval | test | 0.86 | 26.68 | 37.02 | 79.17 | 73.18 | 79.61 | 81.43 | ||
+test | 0.78 | 30.50 | 39.94 | 81.09 | 74.08 | 80.77 | 82.99 | ||
+| RuBQRetrieval | test | 2.54 | 27.09 | 29.74 | 72.27 | 60.34 | 62.11 | 69.12 | ||
+| XQuADRetrieval | validation | 14.28 | 69.50 | 75.24 | 95.26 | 91.97 | 91.53 | 94.15 | ||
+test | -1.07 | -4.64 | 0.97 | -0.20 | 0.60 | 2.66 | -2.92 | ||
+| Average | 4.78 | 32.68 | 38.87 | 66.46 | 60.74 | 63.66 | 64.73 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| RUParaPhraserSTS | test | 40.46 | 51.46 | 49.95 | 74.57 | 69.62 | 71.27 | 74.14 | ||
+| RuSTSBenchmarkSTS | test | 51.02 | 70.40 | 76.15 | 81.54 | 81.52 | 81.36 | 83.67 | ||
+| STS22.v2 | test | 16.95 | 38.21 | 48.97 | 71.00 | 69.81 | 70.75 | 69.92 | ||
+test | 50.63 | 70.13 | 76.28 | 81.53 | 81.53 | 81.36 | 83.44 | ||
+| Average | 39.77 | 57.55 | 62.84 | 77.16 | 75.62 | 76.18 | 77.79 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | ||||||||||
+| ArguAna-PL | test | 7.46 | 10.03 | 22.07 | 38.52 | 43.92 | 46.80 | 40.45 | ||
+| BelebeleRetrieval | test | 30.29 | 35.21 | 53.19 | 92.45 | 89.16 | 89.63 | 90.97 | ||
+| DBPedia-PL | test | 13.01 | 9.41 | 13.34 | 34.88 | 24.07 | 24.83 | 30.45 | ||
+test | 16.36 | 15.28 | 18.87 | 38.80 | 27.67 | 27.98 | 35.09 | ||
+| FiQA-PL | test | 2.47 | 3.10 | 6.56 | 38.85 | 26.67 | 29.02 | 32.71 | ||
+| HotpotQA-PL | test | 19.09 | 12.95 | 16.80 | 60.29 | 49.51 | 48.40 | 55.11 | ||
+test | 24.86 | 19.64 | 23.17 | 62.17 | 51.89 | 51.08 | 56.15 | ||
+| MSMARCO-PL | test | 11.30 | 7.83 | 16.01 | 64.78 | 43.29 | 49.89 | 59.49 | ||
+test | 32.40 | 22.79 | 28.04 | 66.86 | 52.27 | 55.61 | 63.51 | ||
+| NFCorpus-PL | test | 7.40 | 9.99 | 11.26 | 31.39 | 27.00 | 28.71 | 27.66 | ||
+| NQ-PL | test | 3.25 | 4.29 | 7.43 | 54.01 | 29.31 | 30.78 | 43.29 | ||
+| NQ-PLHardNegatives | test | 5.27 | 6.57 | 10.43 | 56.20 | 32.12 | 32.49 | 45.98 | ||
+| Quora-PL | test | 52.30 | 48.28 | 58.88 | 53.57 | 77.74 | 79.20 | 80.65 | ||
+test | 54.25 | 49.15 | 60.60 | 54.82 | 77.87 | 79.01 | 80.57 | ||
+| SCIDOCS-PL | test | 3.14 | 4.38 | 6.60 | 15.38 | 11.96 | 14.35 | 13.73 | ||
+| SciFact-PL | test | 8.82 | 18.42 | 23.21 | 64.88 | 61.69 | 60.32 | 57.68 | ||
+| TRECCOVID-PL | test | 12.15 | 16.74 | 30.85 | 71.73 | 51.25 | 58.40 | 70.81 | ||
+| XPQARetrieval | test | 23.76 | 15.67 | 26.19 | 53.77 | 54.00 | 54.60 | 52.33 | ||
+| Average | 18.20 | 17.21 | 24.08 | 52.96 | 46.19 | 47.84 | 52.04 | |||
+| STS Spearman correlation on cosine similarity | ||||||||||
+| CDSC-R | test | 78.79 | 76.12 | 85.32 | 91.97 | 90.95 | 91.20 | 91.16 | ||
+| SICK-R-PL | test | 53.67 | 48.13 | 53.26 | 72.90 | 74.85 | 75.11 | 78.52 | ||
+| STS22.v2 | test | 20.31 | 16.65 | 30.84 | 49.36 | 45.12 | 50.53 | 45.73 | ||
+test | 57.05 | 49.94 | 61.56 | 82.05 | 80.90 | 81.04 | 82.62 | ||
+| Average | 52.46 | 47.71 | 57.75 | 74.07 | 72.95 | 74.47 | 74.51 | |||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | Languages | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| Retrieval nDCG@10 [%] | |||||||||||
+| BelebeleRetrieval | test | de,en | 66.96 | 52.96 | 69.05 | 92.38 | 92.11 | 92.36 | 92.40 | ||
+| BelebeleRetrieval | test | en,de | 46.64 | 35.59 | 62.72 | 91.80 | 90.18 | 91.19 | 89.84 | ||
+| BelebeleRetrieval | test | fr,en | 79.80 | 64.85 | 73.07 | 92.75 | 88.64 | 90.50 | 92.32 | ||
+| BelebeleRetrieval | test | en,fr | 62.43 | 54.01 | 74.08 | 92.16 | 89.35 | 91.16 | 90.53 | ||
+| BelebeleRetrieval | test | hi,en | 8.07 | 68.55 | 71.82 | 91.48 | 87.95 | 89.73 | 90.46 | ||
+| BelebeleRetrieval | test | en,hi | 2.88 | 65.36 | 72.27 | 87.67 | 82.39 | 85.74 | 83.21 | ||
+| BelebeleRetrieval | test | hi,en | 55.56 | 33.15 | 36.43 | 70.16 | 68.62 | 68.77 | 69.62 | ||
+| BelebeleRetrieval | test | en,hi | 44.96 | 24.18 | 32.34 | 57.53 | 51.68 | 54.83 | 51.97 | ||
+| BelebeleRetrieval | test | jp,en | 25.11 | 67.58 | 71.78 | 91.17 | 88.68 | 90.24 | 91.24 | ||
+| BelebeleRetrieval | test | en,jp | 7.18 | 60.99 | 70.07 | 89.49 | 86.80 | 88.90 | 85.39 | ||
+| BelebeleRetrieval | test | pl,en | 45.49 | 48.50 | 63.12 | 92.22 | 88.98 | 90.17 | 91.70 | ||
+| BelebeleRetrieval | test | en,pl | 28.36 | 36.52 | 58.30 | 90.63 | 89.00 | 90.34 | 88.36 | ||
+| BelebeleRetrieval | test | ru,en | 15.23 | 64.24 | 71.26 | 91.70 | 87.98 | 89.50 | 91.25 | ||
+| BelebeleRetrieval | test | en,ru | 7.67 | 63.76 | 73.50 | 91.82 | 89.75 | 91.20 | 90.16 | ||
+| BelebeleRetrieval | test | es,en | 74.95 | 64.49 | 72.92 | 92.65 | 90.23 | 90.84 | 91.79 | ||
+| BelebeleRetrieval | test | en,es | 58.15 | 53.79 | 72.11 | 91.33 | 89.96 | 91.11 | 90.96 | ||
+| BelebeleRetrieval | test | zh,en | 31.42 | 62.78 | 68.42 | 89.68 | 88.79 | 89.78 | 89.86 | ||
+| BelebeleRetrieval | test | en,zh | 8.69 | 61.78 | 69.85 | 89.43 | 85.59 | 88.81 | 86.22 | ||
+| BelebeleRetrieval | test | zh,en | 27.30 | 62.34 | 67.54 | 89.52 | 87.45 | 88.94 | 89.48 | ||
+| BelebeleRetrieval | test | en,zh | 7.50 | 61.05 | 69.84 | 89.92 | 86.40 | 88.74 | 85.50 | ||
+| BelebeleRetrieval | test | hi,hi | 6.41 | 22.82 | 31.82 | 58.73 | 52.19 | 57.18 | 51.35 | ||
+| BelebeleRetrieval | test | hi,hi | 2.23 | 31.65 | 35.49 | 67.89 | 65.01 | 65.83 | 64.99 | ||
+| CUREv1 | dentistry_and_oral_health | es,en | 9.53 | 11.85 | 22.29 | 44.83 | 37.08 | 40.17 | 47.88 | ||
+| CUREv1 | dermatology | es,en | 21.74 | 12.79 | 18.25 | 50.08 | 36.11 | 41.21 | 56.62 | ||
+| CUREv1 | gastroenterology | es,en | 18.31 | 12.20 | 20.48 | 48.85 | 38.38 | 41.16 | 51.63 | ||
+| CUREv1 | genetics | es,en | 19.82 | 11.87 | 21.38 | 51.52 | 39.55 | 40.01 | 50.42 | ||
+| CUREv1 | neuroscience_and_neurology | es,en | 9.30 | 9.07 | 18.68 | 41.36 | 29.13 | 31.69 | 40.34 | ||
+| CUREv1 | orthopedic_surgery | es,en | 8.69 | 5.04 | 15.17 | 37.66 | 39.86 | 38.10 | 43.00 | ||
+| CUREv1 | otorhinolaryngology | es,en | 8.30 | 6.23 | 14.92 | 38.43 | 32.59 | 33.32 | 44.43 | ||
+| CUREv1 | plastic_surgery | es,en | 11.80 | 8.55 | 18.93 | 42.34 | 34.48 | 36.89 | 45.23 | ||
+| CUREv1 | psychiatry_and_psychology | es,en | 14.34 | 15.93 | 24.81 | 50.45 | 39.39 | 41.25 | 50.48 | ||
+| CUREv1 | pulmonology | es,en | 13.17 | 13.76 | 23.95 | 48.60 | 36.02 | 36.85 | 47.91 | ||
+| CUREv1 | avg | es,en | 13.50 | 10.73 | 19.89 | 45.41 | 36.26 | 38.06 | 47.79 | ||
+| CUREv1 | dentistry_and_oral_health | fr,en | 12.04 | 12.18 | 21.30 | 42.74 | 34.24 | 37.53 | 48.03 | ||
+| CUREv1 | dermatology | fr,en | 27.76 | 15.22 | 19.91 | 50.74 | 34.70 | 38.40 | 56.08 | ||
+| CUREv1 | gastroenterology | fr,en | 21.87 | 12.85 | 19.89 | 48.14 | 34.75 | 37.14 | 50.20 | ||
+| CUREv1 | genetics | fr,en | 26.69 | 13.79 | 22.69 | 49.52 | 33.82 | 34.83 | 47.66 | ||
+| CUREv1 | neuroscience_and_neurology | fr,en | 14.39 | 10.59 | 16.83 | 38.74 | 24.63 | 27.98 | 39.30 | ||
+| CUREv1 | orthopedic_surgery | fr,en | 14.62 | 6.92 | 12.95 | 35.61 | 35.70 | 36.68 | 44.31 | ||
+| CUREv1 | otorhinolaryngology | fr,en | 10.57 | 7.09 | 14.12 | 40.14 | 29.33 | 30.62 | 44.08 | ||
+| CUREv1 | plastic_surgery | fr,en | 20.62 | 9.15 | 19.19 | 42.27 | 33.00 | 34.33 | 45.15 | ||
+| CUREv1 | psychiatry_and_psychology | fr,en | 20.73 | 14.97 | 25.78 | 49.75 | 36.05 | 37.79 | 48.87 | ||
+| CUREv1 | pulmonology | fr,en | 16.56 | 12.42 | 22.73 | 46.90 | 32.95 | 33.96 | 45.83 | ||
+| CUREv1 | avg | fr,en | 18.58 | 11.52 | 19.54 | 44.46 | 32.92 | 34.93 | 46.95 | ||
+test | de,fr | 31.64 | 45.69 | 70.74 | 83.98 | 89.75 | 89.34 | 91.11 | ||
+test | fr,de | 30.14 | 38.42 | 73.18 | 81.67 | 89.41 | 89.00 | 90.83 | ||
+test | de,fr | 37.51 | 53.97 | 72.56 | 81.64 | 89.81 | 88.24 | 89.81 | ||
+test | fr,de | 39.98 | 53.41 | 77.38 | 83.09 | 84.21 | 84.77 | 87.23 | ||
+| MLQARetrieval | test | de,en | 45.85 | 26.14 | 35.89 | 66.97 | 60.34 | 60.86 | 65.22 | ||
+| MLQARetrieval | test | de,es | 34.32 | 27.08 | 39.11 | 70.58 | 65.32 | 65.55 | 69.11 | ||
+| MLQARetrieval | test | de,hi | 2.54 | 31.57 | 42.82 | 68.43 | 62.34 | 63.67 | 65.91 | ||
+| MLQARetrieval | test | de,zh | 6.91 | 26.13 | 35.68 | 66.87 | 59.62 | 62.15 | 64.21 | ||
+| MLQARetrieval | test | en,de | 33.03 | 17.24 | 32.36 | 69.30 | 66.14 | 66.89 | 67.81 | ||
+| MLQARetrieval | test | en,es | 40.98 | 28.19 | 40.57 | 66.38 | 62.79 | 63.06 | 64.67 | ||
+| MLQARetrieval | test | en,hi | 2.10 | 35.97 | 43.03 | 61.04 | 57.01 | 58.68 | 59.43 | ||
+| MLQARetrieval | test | en,zh | 6.43 | 28.74 | 35.76 | 59.75 | 56.45 | 58.25 | 58.11 | ||
+| MLQARetrieval | test | es,de | 32.10 | 21.12 | 36.79 | 73.89 | 68.02 | 67.89 | 71.58 | ||
+| MLQARetrieval | test | es,en | 50.20 | 34.70 | 42.96 | 68.75 | 61.07 | 61.07 | 67.18 | ||
+| MLQARetrieval | test | es,hi | 2.38 | 40.29 | 49.28 | 68.74 | 62.01 | 62.47 | 66.13 | ||
+| MLQARetrieval | test | es,zh | 6.91 | 32.84 | 41.74 | 68.56 | 62.00 | 63.32 | 66.16 | ||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
+| Task | Split | Languages | jc-v1 | nllb-b | nllb-l | je-v3 | jc-v2-s1 | jc-v2-s2 | jc-v2 | ||
+| MLQARetrieval | test | hi,de | 8.05 | 21.71 | 39.25 | 73.69 | 67.16 | 67.60 | 72.32 | ||
+| MLQARetrieval | test | hi,en | 10.28 | 36.62 | 41.89 | 65.76 | 56.77 | 58.33 | 64.18 | ||
+| MLQARetrieval | test | hi,es | 8.33 | 32.66 | 45.79 | 70.49 | 64.31 | 65.37 | 69.20 | ||
+| MLQARetrieval | test | hi,zh | 2.62 | 34.75 | 42.37 | 67.60 | 61.29 | 62.03 | 65.39 | ||
+| MLQARetrieval | test | zh,de | 14.69 | 18.84 | 33.67 | 67.35 | 59.91 | 63.49 | 66.37 | ||
+| MLQARetrieval | test | zh,en | 18.78 | 29.21 | 36.59 | 60.33 | 48.72 | 53.52 | 59.50 | ||
+| MLQARetrieval | test | zh,es | 14.23 | 26.81 | 39.87 | 64.46 | 57.42 | 59.66 | 63.39 | ||
+| MLQARetrieval | test | zh,hi | 2.43 | 36.52 | 44.44 | 61.93 | 54.28 | 57.14 | 60.20 | ||
+| XPQARetrieval | test | en,de | 13.45 | 9.98 | 28.71 | 60.48 | 58.75 | 60.24 | 60.28 | ||
+| XPQARetrieval | test | de,en | 39.28 | 29.44 | 47.12 | 82.12 | 82.80 | 82.95 | 79.45 | ||
+| XPQARetrieval | test | en,es | 20.41 | 14.32 | 30.48 | 53.46 | 49.64 | 52.31 | 53.17 | ||
+| XPQARetrieval | test | es,en | 40.61 | 31.30 | 45.34 | 68.31 | 69.14 | 70.26 | 67.62 | ||
+| XPQARetrieval | test | en,fr | 22.20 | 15.66 | 31.46 | 54.80 | 52.15 | 56.24 | 55.54 | ||
+| XPQARetrieval | test | fr,en | 47.06 | 36.14 | 49.88 | 73.83 | 75.44 | 76.56 | 72.96 | ||
+| XPQARetrieval | test | en,hi | 4.87 | 27.87 | 34.86 | 41.10 | 36.15 | 38.23 | 39.63 | ||
+| XPQARetrieval | test | hi,en | 8.07 | 56.85 | 62.61 | 75.41 | 74.52 | 74.41 | 71.56 | ||
+| XPQARetrieval | test | en,jp | 5.12 | 17.09 | 28.58 | 47.12 | 47.19 | 49.49 | 46.51 | ||
+| XPQARetrieval | test | jp,en | 19.43 | 46.09 | 54.54 | 73.50 | 71.75 | 73.87 | 69.75 | ||
+| XPQARetrieval | test | en,pl | 12.08 | 9.86 | 16.86 | 34.19 | 34.23 | 35.93 | 34.56 | ||
+| XPQARetrieval | test | pl,en | 20.69 | 17.50 | 25.65 | 50.88 | 51.56 | 52.74 | 49.98 | ||
+| XPQARetrieval | test | en,zh | 4.41 | 12.86 | 19.12 | 37.43 | 34.62 | 36.07 | 36.21 | ||
+| XPQARetrieval | test | zh,en | 12.96 | 32.39 | 39.80 | 64.52 | 62.18 | 60.96 | 58.95 | ||
+test | en,ru | 1.04 | 1.55 | 1.24 | 0.22 | 0.83 | -0.20 | -2.13 | ||
+test | en,zh | 0.18 | -1.27 | -1.47 | -0.69 | 3.07 | 2.71 | -0.72 | ||
+| Average | 23.43 | 36.24 | 47.43 | 69.84 | 66.67 | 68.03 | 68.78 | ||||
+| STS Spearman correlation on cosine similarity | |||||||||||
+| IndicCrosslingualSTS | test | en,hi | -11.26 | 58.03 | 58.73 | 67.84 | 62.49 | 61.18 | 72.51 | ||
+| STS17 | test | en,de | 53.67 | 54.88 | 65.87 | 83.66 | 84.18 | 84.03 | 85.89 | ||
+| STS17 | test | es,en | 52.80 | 64.91 | 77.17 | 85.09 | 85.60 | 85.27 | 87.28 | ||
+| STS17 | test | fr,en | 57.46 | 66.04 | 76.21 | 83.77 | 83.51 | 83.47 | 85.91 | ||
+| STS22.v2 | test | de,en | 55.96 | 56.41 | 54.70 | 61.63 | 65.06 | 65.86 | 61.14 | ||
+| STS22.v2 | test | es,en | 69.92 | 56.79 | 65.06 | 81.80 | 81.14 | 82.53 | 81.48 | ||
+| STS22.v2 | test | pl,en | 62.61 | 58.90 | 67.59 | 76.82 | 77.60 | 82.35 | 78.00 | ||
+| STS22.v2 | test | zh,en | 39.82 | 57.88 | 60.61 | 76.33 | 73.93 | 74.16 | 70.47 | ||
+| STS22.v2 | test | de,fr | 39.37 | 42.33 | 44.69 | 55.67 | 53.96 | 60.86 | 60.39 | ||
+| STS22.v2 | test | de,pl | 25.87 | 17.52 | 27.62 | 56.69 | 39.79 | 42.93 | 47.27 | ||
+| STS22.v2 | test | fr,pl | 84.52 | 50.71 | 39.44 | 84.52 | 50.71 | 61.98 | 61.98 | ||
+| Average | 48.25 | 53.13 | 57.97 | 73.98 | 68.91 | 71.33 | 72.03 | ||||
+Evaluated on all tasks included in MTEB version , except the following: MrTidyRetrieval, BrightRetrieval, MSMARCOv2, NeuCLIR2022Retrieval, NeuCLIR2023Retrieval and MIRACLRetrieval.
+These tasks were excluded either due to bugs in the evaluation code or excessive computation times.
 | Dataset - Dimension |
@@
 | Retrieval - nDCG@10 | ||||||||||||
-| Average Retrieval | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
+| Average | 49.33 | 49.32 | 49.19 | 48.67 | 46.37 | 40.66 | ||||||
 | FiQA2018 | 41.93 | 42.04 | 41.79 | 40.96 | 38.75 | 34.27 | ||||||
@@
 | STS - Spearman correlation based on cosine similarity | ||||||||||||
-| Average STS | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
+| Average | 81.29 | 81.27 | 81.26 | 81.24 | 80.78 | 79.56 | ||||||
 | STS12 | 76.72 | 76.72 | 76.85 | 76.97 | 77.13 | 76.44 | ||||||